### 16 Mar 2026 v1

# Mean field equation and Green’s function on the Tate curve

Abstract
In this paper, we study the spectrum of the Laplacian on the Tate curve and construct the associated Green’s function as a finite sum, which can be viewed as the non-Archimedean counterpart of the Green’s function on the flat torus in the Archimedean case. And we establish the existence and uniqueness results of the mean field equation on this space. To address this problem, we begin by presenting the solution structure and proving its uniqueness under specific conditions on the finite quotient. We then establish convergence of the solutions to obtain existence results. Notably, the well-posedness of the solution resembles that in the Archimedean case.
1 Introduction
The mean field equation
$\Delta u+\rho e^{u}=\rho\delta_{0},\ \rho>0$ (1.1)
arises from statistical mechanics and resembles the Liouville equation, the latter stems from the prescribed Gaussian curvature problem but formulated without a point source term on the right-hand side. A more general formulation is given by the Kazdan-Warner equation. When $\rho=4\pi$, (1.1) is related to the Chern-Simons model [4]. On special domains such as $\Omega\subseteq\mathbb{R}^{n}$ or compact manifolds, the existence and uniqueness of the solution has been studied in [5, 21, 22, 6, 7, 17].
The existence of the solution on graphs has also been investigated. Grigor’yan, Lin, and Yang [9] provided necessary conditions for the existence of solutions on finite graphs, while Huang-Lin-Yau [13] established the existence of solutions on finite weighted graphs. Additionally, Ge and Jiang [8] further proved the existence of solutions on infinite graphs.
Ling and Wang [20, 19] studied the mean field equation on the flat torus $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ for $\tau\in\mathbb{C}\backslash\mathbb{R}$. When $\rho\neq 8m\pi$ with $m\in\mathbb{Z}$, there always exist a solution as the Leray-Schauder degree is nonzero, independent of the torus shape. When $\rho=8\pi$, they demonstrated that the existence and uniqueness of the solution is equivalent to the number of critical points of the Green’s function
$\Delta G(z,w)=\delta_{w}(z)-\frac{1}{V}$ (1.2)
where $V$ denotes the volume of the torus. Specifically, if the mean field equation admits a unique solution, then the Green’s function possesses five critical points, otherwise it has only three.
Since these Liouville-type equations have numerous applications in quantum field theory and string theory [23, 10, 1], and the Green’s function underlies the computation of Veneziano amplitudes. It is natural to investigate a $p$-adic analogue of this problem. Analyses of $p$-adic pseudo-differential operators can be found in [24].
In the genus zero case, Huang-Stoica-Yau-Zhong [15, 16] defined the $p$-adic version of the flat Laplacian via regularizing Vladimirov derivation as
$D^{s}u(x)=\int_{\mathbb{Q}_{p}}\frac{u(z)-u(x)}{|z-x|_{p}^{1+s}}dz$
and constructed the associated Green’s function, where $dz$ is the Haar measure on $\mathbb{Q}_{p}$. Besides, in the upper-half plane, Hassan [11] constructed the higher Green’s function.
For genus one, Huang-Rohrlich-Sun-Whyman [14] defined the $p$-adic version of the flat Laplacian on the Tate curve as
$Du(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}dz$
where $|q|_{p}<1$. It is easy to check that the operator $D$ is self-adjoint, negative semi-definite and preserves locally constant functions under the inner product induced by the Haar measure $\frac{dx}{|x|_{p}}$ on $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$. They proved the Green’s function only depends on $d(x,y)=\frac{|x-y|_{p}}{\text{max}\{|x|_{p},|y|_{p}\}}$ and provided a series expression for the Green’s function by correcting the logarithmic singularity on the diagonal.
Remark 1.1.
$\mathbb{Q}_{p}/(\mathbb{Z}+\mathbb{Z}\tau)$ is not a suitable object because $\mathbb{Z}+\mathbb{Z}\tau$ is not closed in $\mathbb{Q}_{p}$, hence the quotient topology is not Hausdorff. In the Archimedean case, we consider the exponential map: $x\mapsto e^{2\pi ix}$, yielding $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)\rightarrow\mathbb{C}^{\times}/(e^{2\pi i\tau})^{\mathbb{Z}}$. So the Tate curve $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ with $|q|_{p}<1$ can be viewed as the non-Archimedean analogy of the flat torus.
Remark 1.2.
Since multiplication by a unit in $\mathbb{Z}_{p}^{*}$ leaves $D$ invariant, we may restrict to the case $q=p^{m}$, then we choose the fundamental domain as $\cup_{k=0}^{m-1}p^{l}\mathbb{Z}_{p}^{*}$. One can verify that the definition is independent of the choice of the fundamental domain.
For more general cases, Bradley derived the Green’s function via heat kernel on a compact domain [2] and showed the existence of the heat kernel and the Green’s function for fractional Laplacian with $s>1$ on compact manifolds [3].
In this paper, we focus on the mean field equation and the Green’s function on the Tate curve as
$DG(x,y)=\delta_{y}(x)-\frac{1}{V}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ (1.3)
$Du+\rho e^{u}=\rho\delta_{y}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ (1.4)
A standard approach to constructing the Green’s function employs the spectrum of the operator $D$ and its eigenfunctions. When $m=1$, the fundamental domain is a locally compact group, allowing us to obtain its spectrum and eigenfunctions via characters, some of these works have been established in [12]. We extend the results to the general $m$ case through zero extension although the operator $D$ is a nonlocal operator, and make minor modifications in the case when the eigenvalue is zero. Using orthogonality of the characters, we express the Green’s function as a finite sum, the resulting formula differs form that in [14].
Let $\boldsymbol{L}=\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ and $\boldsymbol{R}=-(1-p^{-1})\boldsymbol{L}$, where $(\boldsymbol{P})_{ij}=p^{-|i-j|}$, so we have orthogonal matrix $\boldsymbol{Q}$ to diagonalize $\boldsymbol{R}$. And we define $\phi_{h,c,n}(x)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\{\frac{c\text{ln}\frac{x}{a_{0}}}{p^{n}}\}_{p}}$, where $x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, $0\leq h\leq p-2$ and $c\in\mathbb{Z}_{p}^{*}$. Then our first main result is:
Theorem 1.3 (The spectrum of $D$ and the corresponding Green’s function).
The eigenvalues of $D$ for $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ are $\lambda_{0,l}\in\sigma(\boldsymbol{R})$ with $\lambda_{0,0}=0$, and $\lambda_{n,l}=-p^{n-1}-p^{n-2}+\frac{p^{-l}+p^{l+1-m}}{p}$ for $n\geq 1$, with the corresponding eigenfunctions $\phi_{0}^{l}(x)=\sum_{k=0}^{m-1}q_{k,l}\boldsymbol{1}_{p^{k}\mathbb{Z}_{p}^{*}}(x)$ and $\phi_{h,c,n}^{l}(x)=\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$, where $l\in\{0,1,2,\dots,m-1\}$. And the multiplicities are $m_{0,0}=1,m_{1,l}=p-2$ and $m_{n,l}=(p-1)^{2}p^{n-2}$ for $n\geq 2$.
And up to a constant $C(|y|_{p})$, the corresponding Green’s function (1.3) is
$G(x,y)=\frac{p^{M+1}}{p^{M+1}+p^{M}-(p^{-l}+p^{l+1-m})}-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-(p^{-l}+p^{l+1-m})}+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{(1-p^{-1})\lambda_{0,k}}$
when $v_{p}(x)=v_{p}(y)=l$ and $v_{p}(d(x,y))=M$.
Inspired by the existence results for the Green’s function [14], we first consider the mean field equation on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$. Since there are finitely many points in the finite quotient, we can solve the equation by numerical experiments. Based on observations from numerical results, we investigate a universal upper bound of the solution and establish uniqueness when the parameter $\rho\leq C(p,m)$. We then study the convergence of the solutions to obtain existence on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$. Similar to the arguments on the finite quotient, the uniqueness can be preserved, yielding a unique solution. The optimal upper bound of $\rho$ can be viewed as analogous to the $8\pi$ threshold in Archimedean case. Then our second main result is:
Theorem 1.4.
When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then there exists a unique solution of the mean field equation (1.4). Where $\lambda_{1}$ is the smallest nonzero eigenvalue of $\boldsymbol{L}$.
Our ultimate goal is to connect the Green’s function with the mean field equation to explore the critical point properties of the Green’s function. Since the normal derivative on the Green’s function is trivial due to its local constancy, our first step will be to define a suitable derivative on the Green’s function in future work.
1.1 Outline of the paper
In section 2, we study the spectrum of $D$ for the case $m=1$ by using character theory and construct the Green’s function and heat kernel as a finite sum, and we promote the result to the general $m$ case.
In section 3, we study the mean field equation on finite quotients of the Tate curve. We derive a uniform positive upper bound $C_{p}$(depending only on $p$) for the solution and demonstrate its central symmetry, cyclic symmetry, and monotonicity. Furthermore, we prove that the solution is radial symmetric and dependents only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$ when $\rho\leq C_{p}(1-p^{-m})$, which is similar to the property of the Green’s function. Lastly we establish uniqueness for $\rho\leq C_{p}\mu_{1}$, where $\mu_{1}$ is the smallest nonzero eigenvalue of D-D.
In section 4, we show that the solutions on the finite quotient converge to a solution on the Tate curve while preserving their structural properties, thereby establishing existence, and we prove that under the same condition the solution is unique.
2 Spectrum of D and the Green’s function
This section focuses on the spectrum of $D$ and its eigenfunctions. First, we consider the case $m=1$, then the fundamental domain is $\mathbb{Z}_{p}^{*}$, which is a compact multiplicative group. We analyze $D$ via characters of $\mathbb{Z}_{p}^{*}$. After that, we construct the Green’s function with a finite sum. Lastly, we generalize it to the general $m$ case.
2.1 Multiplicative Characters of $\mathbb{Z}_{p}^{*}$
Since $\mathbb{Z}_{p}^{*}\cong(\mathbb{Z}/p\mathbb{Z})^{\times}\times(1+p\mathbb{Z}_{p})$, where the first factor is cyclic and the second is torsion free, it suffices to determine the characters of both components.
Let $g$ be a primitive root of $(\mathbb{Z}/p\mathbb{Z})^{\times}$, as the dual group of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ is isomorphic to itself, so for $m,h\in\{0,1,\dots,p-2\}$, the characters for $(\mathbb{Z}/p\mathbb{Z})^{\times}$ are given by
$\chi_{h}(g^{m})=e^{\frac{2\pi ih}{p-1}m}$
As the dual group of $\mathbb{Z}_{p}$ is isomorphic to $\mathbb{Q}_{p}/\mathbb{Z}_{p}$, so the characters on $\mathbb{Z}_{p}$ are $\chi_{t}(x)=e^{2\pi i\{tx\}_{p}}$, where $x\in\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. And we notice that there is a continuous isomorphism $\mathbb{Z}_{p}\cong 1+p\mathbb{Z}_{p},x\mapsto e^{px}$. So the characters of $1+p\mathbb{Z}_{p}$ can be pulled back from the isomorphism as
$\chi_{t}(x)=e^{2\pi i\{t\frac{\text{ln}x}{p}\}_{p}}$
where $x\in 1+p\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. Denoting $t=a_{-n}p^{-n}+a_{-n+1}p^{-n+1}+\dots+a_{-1}p^{-1}=\frac{a_{-n}+a_{-n+1}p+\dots+a_{-1}p^{n-1}}{p^{n}}$, and taking $c=a_{-n}+a_{-n+1}p+\dots+a_{-1}p^{n-1}$, we rewrite the character $\chi_{t}(x)$ as
$\chi_{c,n+1}(x)=e^{2\pi i\{\frac{c\text{ln}x}{p^{n+1}}\}_{p}}$
So $\forall x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, the character of $\mathbb{Z}_{p}^{*}$ is trivial character $1$ and
$\chi_{h,c,n}(x)=\chi_{h}(a_{0})\chi_{c,n}(\frac{x}{a_{0}})=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\{\frac{c\text{ln}\frac{x}{a_{0}}}{p^{n}}\}_{p}}$ (2.1)
where $h\in\{0,1,\dots,p-2\},n\in\mathbb{N_{+}}$. And the conductor of $\chi_{h,c,n}$ is $n$.
2.2 Eigenvalues and eigenfunctions for $D$ when $m=1$
Lemma 2.1.
The characters $\{\chi_{h,c,n}\}_{h,c,n}$ and the trivial character $1$ are orthogonal.
Proof.
First, $\langle 1,1\rangle=1-p^{-1}$ and $\langle\chi_{h,c,n},1\rangle=0$ as $\chi_{h,c,n}$ is nontrivial on $\mathbb{Z}_{p}^{*}$. Next,
$\langle\chi_{h,c,n},\chi_{h^{\prime},c^{\prime},n^{\prime}}\rangle$ $=\int_{\mathbb{Z}_{p}^{*}}\chi_{h,c,n}(x)\overline{\chi_{h^{\prime},c^{\prime},n^{\prime}}(x)}dx$
$=\int_{\mathbb{Z}_{p}^{*}}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})\text{ln}\frac{x}{a_{0}}\}_{p}}dx$
$=\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{1+p\mathbb{Z}_{p}}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})\text{ln}\frac{x}{a_{0}}\}_{p}}dx$
$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{\mathbb{Z}_{p}}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})y\}_{p}}dy$
$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\boldsymbol{1}_{\mathbb{Z}_{p}}(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})$
$=(1-p^{-1})\delta_{h,c,n}^{h^{\prime},c^{\prime},n^{\prime}}$
thus establishing orthogonality. ∎
Theorem 2.2.
The eigenvalues of $D$ are $\lambda_{0}=0$ and $\lambda_{n}=-p^{n-1}-p^{n-2}+\frac{2}{p}$ for $n\geq 1$, with the corresponding eigenfunctions $\phi_{0}=1$ and $\phi_{h,c,n}=\chi_{h,c,n}$. The multiplicities are $m_{0}=1,m_{1}=p-2$ and $m_{n}=(p-1)^{2}p^{n-2}$ for $n\geq 2$.
Proof.
Since the conductor of $\chi_{h,c,n}$ is $n$, it is trivial on $1+p^{n}\mathbb{Z}_{p}$ but nontrivial on $1+p^{n-1}\mathbb{Z}_{p}$. By the first orthogonality relation of the character, $\forall s\in\{1,2,\dots,n-1\}$, we have
$\sum_{x\in(1+p^{s}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})}\chi_{h,c,n}(x)=0$
so we only need to consider $z=a_{0}+a_{1}p+\dots+a_{n-1}p^{n-1}$ and $x=b_{0}+b_{1}p+\dots+b_{n-1}p^{n-1}$. Then we have
$D\chi_{h,c,n}(x)$ $=\int_{\mathbb{Z}_{p}^{*}}\frac{\chi_{h,c,n}(z)-\chi_{h,c,n}(x)}{|z-x|_{p}^{2}}dz$
$=\mu(1+p^{n}\mathbb{Z}_{p})\int_{\mathbb{Z}_{p}^{*}/(1+p^{n}\mathbb{Z}_{p})}\frac{\chi_{h,c,n}(z)-\chi_{h,c,n}(x)}{|z-x|_{p}^{2}}d\delta$
$=p^{-n}\sum_{j=0}^{n-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{n-1}\end{subarray}}p^{2j}(\chi_{h,c,n}(z)-\chi_{h,c,n}(x))$
$=-p^{-n}((p-2)p^{n-1}+\sum_{d=1}^{n-2}(p-1)p^{n-1-d}p^{2d}+(p-1-(-1))p^{2(n-1)})\chi_{h,c,n}(x)$
$=-p^{-n}((p-2)p^{n-1}+\sum_{d=1}^{n-2}(p-1)p^{n-1+d}+p^{2n-1})\chi_{h,c,n}(x)$
$=-p^{-n}(p^{2n-1}+p^{2n-2}-2p^{n-1})\chi_{h,c,n}(x)$
$=-(p^{n-1}+p^{n-2}-\frac{2}{p})\chi_{h,c,n}(x)$
Hence $\chi_{h,c,n}$ is an eigenfunction of $D$ with eigenvalue $p^{n-1}-p^{n-2}+\frac{2}{p}$ for $n\geq 1$.
Next, for any locally constant eigenfunction $f$ in $\mathbb{Z}_{p}^{*}$, we have the Fourier transform
$\hat{f}(\xi)=\int_{\mathbb{Z}_{p}^{*}}f(x)\overline{\chi_{\xi}(x)}dx$
then
$f(x)=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x)$
so we have
$Df(x)$ $=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)(D\chi_{\xi}(x))$
$=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\lambda_{\xi}\chi_{\xi}(x)$
$=\lambda_{f}\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x)$
thus
$\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}(\lambda_{\xi}-\lambda_{f})\hat{f}(\xi)\chi_{\xi}(x)=0$
By orthogonality of eigenfunctions, we have $\lambda_{\xi}=\lambda_{f}$ if $\lambda_{\xi}$ belongs to the expansion of $f$. That means $\text{Char}(\mathbb{Z}_{p}^{*})$ is a complete eigenbasis of all eigenfunctions of $D$.
Lastly, the multiplicity of eigenvalues comes from counting the number of $h$ and $c$. ∎
2.3 Construction of the Green’s function when $m=1$
Using the spectrum of $D$, we construct the Green’s function as
$G(x,y)=\frac{1}{1-p^{-1}}\sum_{\phi_{h,c,n}\in\text{Char}(\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{h,c,n}(x)\overline{\phi_{h,c,n}(y)}}{\lambda_{n}}+C$ (2.2)
Let $H_{n}$ denote the set of characters with conductor at most $n$ and define $C_{n}:=H_{n}\backslash H_{n-1}$. By the second orthogonality relation of the character, we have
$\sum_{\phi\in H_{n}}\phi(x)\overline{\phi(y)}=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})|H_{n}|$ (2.3)
where $|H_{n}|=(p-1)p^{n-1}=\varphi(p^{n})$. And we notice that if $v_{p}(x-y)=M$ and $n\geq M+1$, the sum (2.3) vanishes. [^1] The idea of the finite sum formula for the case $m=1$ originated from Ekam’s independent study. Therefore,
$G(x,y)$ $=\frac{1}{1-p^{-1}}\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n}}+C$ (2.4)
$=\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n}}-\frac{\varphi(p^{M})}{\lambda_{M+1}})+C$
$=-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-2}+\frac{p^{M+1}}{p^{M+1}+p^{M}-2}+C$
which is expressed as a finite sum.
2.4 General $m$ case
For general $m$, the fundamental domain is $\cup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}$. First, we consider $\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$ for $n\geq 1$. We have
$D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)$ (2.5)
$=$ $|x|_{p}\int_{\cup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$
$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz-\left(|x|_{p}\int_{\cup_{k\neq l}p^{k}\mathbb{Z}_{p}^{*}}\frac{1}{|z-x|_{p}^{2}}dz\right)\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
where the first term is
$|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$ (2.6)
$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{p^{-2l}|p^{-l}z-p^{-l}x|_{p}^{2}}dz$
$=$ $|p^{-l}x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|p^{-l}z-p^{-l}x|_{p}^{2}}dp^{-l}z$
$=$ $|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz$
when $v_{p}(x)=l$, (LABEL:S:eq_3) reduces to
$\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)}{|z-p^{-l}x|_{p}^{2}}dz=(D_{0}\phi_{h,c,n})(p^{-l}x)=\lambda_{n}\phi_{h,c,n}(p^{-l}x)$
while the second term is
$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}}{\text{max}\{p^{-k},p^{-l}\}^{2}}p^{-k}(1-p^{-1})\phi_{h,c,n}(p^{-l}x)$
Then the coefficient of $\phi_{h,c,n}(p^{-l}x)$ simplifies to
$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}p^{-k}}{\text{max}\{p^{-k},p^{-l}\}^{2}}(1-p^{-1})$
$=$ $-(1-p^{-1})(\sum_{k=0}^{l-1}p^{-l}p^{-k}p^{2k}+\sum_{k=l+1}^{m-1}p^{-l}p^{-k}p^{2l})$
$=$ $-(1-p^{-1})(p^{-l}\sum_{k=0}^{l-1}p^{k}+\sum_{k=l+1}^{m-1}p^{-(k-l)})$
$=$ $\frac{p^{-l}+p^{l+1-m}-2}{p}$
thus (LABEL:S:eq_2) becomes
$-(p^{n-1}+p^{n-2}-\frac{2}{p}-\frac{p^{-l}+p^{l+1-m}-2}{p})\phi_{h,c,n}(p^{-l}x)$
$=$ $-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})\phi_{h,c,n}(p^{-l}x)$
when $v_{p}(x)\neq l$, (LABEL:S:eq_3) becomes
$\frac{|p^{-l}x|_{p}}{\text{max}\{1,|p^{-l}x|_{p}\}^{2}}\int_{\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(z)dz=0$
since $\phi_{h,c,n}$ is orthogonal to $\phi_{0}=1$ by Lemma 2.1, and the second term vanishes. Denoting $\lambda_{n,l}=-p^{n-1}-p^{n-2}+\frac{p^{-l}+p^{l+1-m}}{p}$, we obtain
$D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)=\lambda_{n,l}\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
which means $\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$ is the eigenfunction of $D$ with the eigenvalue $\lambda_{n,l}$.
When $n=0$, we consider $\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$, similarly, we have
$D\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{1-\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz+\frac{p^{-l}+p^{l+1-m}-2}{p}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
so when $v_{p}(x)=l$, the first part vanishes, and when $v_{p}(x)\neq l$, the first part becomes
$\frac{|p^{-l}x|_{p}(1-p^{-1})}{\text{max}\{1,|p^{-l}x|_{p}\}^{2}}=(1-p^{-1})p^{-|v_{p}(x)-l|}$
denoting this vector by $\boldsymbol{r}_{l}=\begin{pmatrix}r_{0l}\\ r_{1l}\\ \vdots\\ r_{m-1,l}\end{pmatrix}$. We want to use the linear combination of them to construct the eigenvalue and eigenfunction:
$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\sum_{l=0}^{m-1}c_{l}\boldsymbol{r}_{l}=(\boldsymbol{r}_{0},\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{m-1})\begin{pmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{m-1}\end{pmatrix}=\boldsymbol{R}\boldsymbol{c}$
where we require that
$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\boldsymbol{c}$
so we only need to solve
$\boldsymbol{R}\boldsymbol{c}=\lambda\boldsymbol{c}$
[…]
Lemma 2.3.
The sets $\{\phi_{0}^{l}\}_{l},\{\phi_{h,c,n}^{l}\}_{h,c,n,l}$ form a complete orthogonal eigenbasis for $D$.
Proof.
The orthogonality of $\{\phi_{0}^{l}\}_{l}$ follows from the symmetry of $\boldsymbol{R}$. For $\phi_{h,c,n}^{l}$, by Lemma 2.1 we have
$\langle\phi_{h,c,n}^{l},\phi_{h^{\prime},c^{\prime},n^{\prime}}^{l^{\prime}}\rangle$ $=\delta_{l}^{l^{\prime}}\langle\phi_{h,c,n}^{l},\phi_{h^{\prime},c^{\prime},n^{\prime}}^{l}\rangle$
$=\delta_{l}^{l^{\prime}}\int_{p^{l}\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(p^{-l}x)\overline{\phi_{h,c,n}(p^{-l}x)}\frac{dx}{|x|_{p}}$
$=\delta_{l}^{l^{\prime}}\int_{\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(x)\overline{\phi_{h,c,n}(x)}dx$
$=\delta_{l}^{l^{\prime}}\langle\phi_{h,c,n},\phi_{h^{\prime},c^{\prime},n^{\prime}}\rangle$
$=(1-p^{-1})\delta_{h,c,n,l}^{h^{\prime},c^{\prime},n^{\prime},l^{\prime}}$
which is also orthogonal. Lastly, $\phi_{h,c,n}^{l},\phi_{0}^{l^{\prime}}>=0$ is trivial as $\phi_{h,c,n}^{l}$ is not a constant function.
For any eigenfunction $f$, we have the decomposition $f=f_{c}+f_{0}+f_{1}+\dots+f_{m-1}$, where $f_{c}$ is constant on each $p^{k}\mathbb{Z}_{p}^{*}$ and $\text{supp}f_{l}\subseteq p^{l}\mathbb{Z}_{p}^{*}$ with zero DC component (i.e. the Fourier expansion of $f_{l}$ don’t have trivial character). We can define the Fourier transform of $f_{l}$ by pulling it back to $\mathbb{Z}_{p}^{*}$ as $\widehat{f_{l}}(\xi):=\widehat{f_{l}(p^{-l}x)}(\xi)$. Then we can expand $f$ by Fourier series, and the characters are exactly the eigenfunctions. Similar to the discussion of $m=1$, this establishes completeness. ∎
Thus we have proved:
Theorem 2.4.
The eigenvalues of $D$ are $\lambda_{0,l}\in\sigma(\boldsymbol{R})$ with $\lambda_{0,0}=0$, and $\lambda_{n,l}=-p^{n-1}-p^{n-2}+\frac{p^{-l}+p^{l+1-m}}{p}$ for $n\geq 1$, with the corresponding eigenfunctions $\phi_{0}^{l}(x)=\sum_{k=0}^{m-1}q_{k,l}\boldsymbol{1}_{p^{k}\mathbb{Z}_{p}^{*}}(x)$ and $\phi_{h,c,n}^{l}(x)=\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$, where $l\in\{0,1,2,\dots,m-1\}$. And the multiplicities are $m_{0,0}=1,m_{1,l}=p-2$ and $m_{n,l}=(p-1)^{2}p^{n-2}$ for $n\geq 2$.
Remark 2.5.
$\lambda_{n,l}$ may not be disjoint as $\lambda_{n,l}=\lambda_{n,m-l}$ for $0\leq l\leq m-1,n\geq 1$.
Then we have the Green’s function as
$G(x,y)=\frac{1}{1-p^{-1}}(\sum_{k=0}^{m-1}\sum_{\phi_{n}^{k}\in\text{Char}(p^{k}\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{n}^{k}(x)\overline{\phi_{n}^{k}(y)}}{\lambda_{n,k}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}})+C(|y|_{p})$ (2.7)
Similarly, let $H_{n}^{l}$ denote the set of characters generated from $p^{l}\mathbb{Z}_{p}^{*}$ with conductor at most $n$ and define $C_{n}^{l}:=H_{n}^{l}\backslash H_{n-1}^{l}$. Consider $\sum_{k=0}^{m-1}\displaystyle\sum_{\phi^{k}\in H_{n}^{k}}\phi^{k}(x)\overline{\phi^{k}(y)}$. When $v_{p}(x)\neq v_{p}(y)$ (i.e. $v_{p}(d(x,y))=0$), the sum vanishes. When $v_{p}(x)=v_{p}(y)=l$ (i.e. $v_{p}(d(x,y))>0$), then
$\sum_{\phi^{l}\in H_{n}^{l}}\phi^{l}(x)\overline{\phi^{l}(y)}$ $=\sum_{\phi^{0}\in H_{n}^{0}}\phi^{0}(p^{-l}x)\overline{\phi^{0}(p^{-l}y)}$
$=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})\varphi(p^{n})$
So when $v_{p}(d(x,y))=M$ and $v_{p}(x)=v_{p}(y)=l$, we have the Green’s function
$G(x,y)=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n,l}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}})+C(|y|_{p})$ (2.8)
$=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n,l}}-\frac{\varphi(p^{M})}{\lambda_{M+1,l}}+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{\lambda_{0,k}})+C(|y|_{p})$
$=$ $-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-(p^{-l}+p^{l+1-m})}+\frac{p^{M+1}}{p^{M+1}+p^{M}-(p^{-l}+p^{l+1-m})}$
$+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{(1-p^{-1})\lambda_{0,k}}+C(|y|_{p})$
Finally, we verify that $G(x,y)$ is actually the Green’s function:
When $x\neq y$,
[…]
$=\frac{1}{1-p^{-1}}((\sum_{n=1}^{M}(\varphi(p^{n})-\varphi(p^{n-1}))-\varphi(p^{M}))\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=1}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k})$
$=\frac{1}{1-p^{-1}}(-\varphi(1)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=0}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}-q_{v_{p}(x)0}q_{v_{p}(y)0})$
$=\frac{1}{1-p^{-1}}(-\delta_{v_{p}(x)}^{v_{p}(y)}+\delta_{v_{p}(x)}^{v_{p}(y)}-\frac{1}{m})$
$=-\frac{1}{m(1-p^{-1})}$
as $\boldsymbol{Q}$ is an orthogonal matrix and $\boldsymbol{q}_{0}=\frac{1}{\sqrt{m}}\boldsymbol{1}$.
Then $DG(x,y)+\frac{1}{m(1-p^{-1})}$ is a distribution supported on the diagonal. And we notice that
$\int_{\cup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}DG(x,y)+\frac{1}{m(1-p^{-1})}dx$ (2.10)
[…]
$=$ $1$
So $G(x,y)$ satisfies $DG(x,y)=\delta_{y}(x)-\frac{1}{m(1-p^{-1})}$.
We can compute the heat kernel by using spectrum of $D$ as
$H(t,x,y)=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}e^{\lambda_{n,l}t}\phi(x)\overline{\phi(y)}+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)})$ (2.11)
$=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}e^{\lambda_{n,l}t}(\varphi(p^{n})-\varphi(p^{n-1}))-e^{\lambda_{M+1,l}t}\varphi(p^{M})+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k})$
$=$ $(p-1)\sum_{n=1}^{M}e^{-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})t}p^{n-1}-e^{-(p^{M}+p^{M-1}-\frac{p^{-l}+p^{l+1-m}}{p})t}p^{M}$
$+\frac{p}{p-1}\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k}$
which is also an important fundamental quantity.
3 Mean field equation on the finite quotient
In this section, we consider the mean field equation on the Tate curve
$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ (3.1)
To study this equation, we first consider this equation on a finite quotient
$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ (3.2)
i.e. by truncating at the $d$-th $p$-adic digits. Points are of the form $x=p^{l}(b_{0}+b_{1}p+\dots+b_{d-1}p^{d-1})$ with $l\in\{0,1,\dots,m-1\}$, $b_{0}\in\{1,2,\dots,p-1\}$ and $b_{i}\in\{0,1,\dots,p-1\}$ for $1\leq i\leq d-1$. As the haar measure on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ is $\mu^{\times}(p^{k}(1+p^{d}\mathbb{Z}_{p}))d\delta=p^{-d}d\delta$, where $d\delta$ is the counting measure, denote $z=p^{k}(a_{0}+a_{1}p+\dots+a_{d-1}p^{d-1})$, we have
$Du(x)$ $=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})}|z|_{p}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}p^{-d}d\delta$ (3.3)
$=p^{-d}\sum_{\begin{subarray}{c}0\leq k\leq m-1\\ 1\leq a_{0}\leq p-1\\ 0\leq a_{i}\leq p-1\\ 1\leq i\leq d-1\end{subarray}}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x))$
So the flat Laplacian on the quotient space of the Tate curve reduces to a graph Laplacian on a finite weighted graph. The existence of this solution is guaranteed based on [13]. Let $\boldsymbol{u}=u(p^{l}\sum_{i=0}^{d-1}b_{i}p^{i})$, $e^{\boldsymbol{u}}=e^{u(p^{l}\sum_{i=0}^{d-1}b_{i}p^{i})}$, and define $\boldsymbol{A}=p^{-d}\boldsymbol{B}$ as the coefficient matrix of $\boldsymbol{u}$. Note that $\delta_{y}$ becomes $p^{d}\boldsymbol{e}_{y}$ owing to the haar measure is $p^{-d}d\delta$. Then (3.2) can be rewritten as a matrix equation
$\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\rho p^{d}\boldsymbol{e}_{y}$ (3.4)
From (3.3), $\boldsymbol{A}\in\mathbb{R}^{m(p-1)p^{d-1}\times m(p-1)p^{d-1}}$ is a symmetric matrix with zero row and colum sums. Moreover, $\boldsymbol{A}$ is a block matrix where each block is a symmetric circulant matrix, i.e. $\boldsymbol{A}=(\boldsymbol{A}_{ij})$, $1\leq i,j\leq m$ with $\boldsymbol{B}_{ij}=p^{-|i-j|}\boldsymbol{1}\boldsymbol{1}^{\top}$ for $i\neq j$.
Remark 3.1.
Since the finite quotient can converge to the Tate curve in the sence of Gromov-Hausdorff distance as $d\rightarrow\infty$, the graph Laplacian on this finite graph can converge to the flat Laplacian on the Tate curve. Thus (3.2) serves as a discrete approximation of (3.1).
3.1 Structure of the solution
In this part, we present structural properties of the solution to (3.4). These properties are observed from numerical results.
Lemma 3.2.
If $\boldsymbol{u}$ solves (3.4), then $u_{i}\leq d\text{ln}p,\forall i$. The equality holds if and only if $p=2,m=1$.
Proof.
From (3.4), we have
$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=\rho p^{d}\delta_{iy},\ \forall i$ (3.5)
Summing over $i$ yields
$\sum_{i,j}a_{ij}u_{j}+\rho\sum_{i}e^{u_{i}}=\rho p^{d}$
since $\sum_{i,j}a_{ij}u_{j}=\sum_{j}u_{j}\sum_{i}a_{ij}=0$, we obtain
$\sum_{i}e^{u_{i}}=p^{d}$
then $e^{u_{i}}\leq p^{d}$, thus $u_{i}\leq d\text{ln}p$ for all $i$. And the equality holds if and only if $p=2,m=1$ as there is one term in the sum. ∎
Lemma 3.3.
$u_{y}$ is the only smallest in $\boldsymbol{u}$.
Proof.
Let $m=\text{min}_{i}\ u_{i}$. Suppose $\exists\ i\neq y$, s.t. $u_{i}=m$. From (3.5) with $i\neq y$,
$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=0$
then $\rho e^{u_{i}}+a_{ii}u_{i}=-\sum_{j\neq i}a_{ij}u_{j}\leq-m\sum_{j\neq i}a_{ij}=ma_{ii}$, so we have
$\rho e^{u_{i}}\leq a_{ii}(u_{i}-m)=0$
a contradiction. ∎
Lemma 3.4.
If $\boldsymbol{u}$ solves $\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\boldsymbol{b}$, then $\boldsymbol{J}\boldsymbol{u}$ solves $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\boldsymbol{J}\boldsymbol{b}$, where $\boldsymbol{J}=\begin{pmatrix}&&&1\\ &&1&\\ &\iddots&&\\ 1&&&\end{pmatrix}$ is the anti-diagonal permutation matrix.
Proof.
As $\boldsymbol{A}$ is a symmetric circulant matrix, then it is a centrosymmetric matrix, so we have $\boldsymbol{A}\boldsymbol{J}=\boldsymbol{J}\boldsymbol{A}$. As $\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\boldsymbol{b}$, then $\boldsymbol{J}\boldsymbol{A}\boldsymbol{u}+\rho\boldsymbol{J}e^{\boldsymbol{u}}=\boldsymbol{J}\boldsymbol{b}$, which means $\boldsymbol{A}\boldsymbol{J}\boldsymbol{u}+\rho e^{\boldsymbol{J}\boldsymbol{u}}=\boldsymbol{J}\boldsymbol{b}$, thus $\boldsymbol{J}\boldsymbol{u}$ is a solution of $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\boldsymbol{J}\boldsymbol{b}$. ∎
Let $\boldsymbol{u}=\begin{pmatrix}\boldsymbol{u}_{1}\\ \boldsymbol{u}_{2}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, $\boldsymbol{\sigma}=\begin{pmatrix}&&&1\\ 1&&&\\ &\ddots&&\\ &&1&\end{pmatrix}\in\mathbb{R}^{(p-1)p^{d-1}\times(p-1)p^{d-1}}$ be the shift matrix on blocks of size $(p-1)p^{d-1}$, and define $\sigma_{k}(\boldsymbol{u})=\begin{pmatrix}\boldsymbol{u}_{1}\\ \vdots\\ \boldsymbol{\sigma}\boldsymbol{u}_{k}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}$. When $i\neq j$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ as $\boldsymbol{B}_{ij}=p^{-|i-j|}\boldsymbol{1}\boldsymbol{1}^{\top}$ when $i\neq j$.
Lemma 3.5.
If $\boldsymbol{u}$ solves $\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\boldsymbol{b}$, then $\sigma_{k}(\boldsymbol{u})$ solves $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\sigma_{k}(\boldsymbol{b})$.
Proof.
As $\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\boldsymbol{b}$, so we have
$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{k}}=\boldsymbol{b}_{k}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
multiplying the $k$-th block by $\boldsymbol{\sigma}$ gives
$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{\sigma}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho\boldsymbol{\sigma}e^{\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
which is equivalent to
$\begin{cases}\displaystyle\sum_{j\neq k}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\boldsymbol{A}_{1k}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\boldsymbol{A}_{kk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{\sigma}\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k}\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\boldsymbol{A}_{mk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
Thus $\sigma_{k}(\boldsymbol{u})$ is a solution of $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\sigma_{k}(\boldsymbol{b})$. ∎
Lemma 3.4 and Lemma 3.5 imply certain symmetric properties: given one solution, we can construct many solutions by folding or rotating the solution. However, if the solution is radial symmetric, then Lemma 3.4 and Lemma 3.5 are trivial. From numerical results we can see that there indeed have radial symmetric solutions, and by the following theorem we proved that there only exist radial symmetric solutions when $\rho$ is small.
Theorem 3.6.
When $\rho\leq p^{-d}-p^{-d-m}$, denote $u_{y}(x)$ be the solution of (3.4), then $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$. Moreover, if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ and $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$, which means that $u(x)$ is radial symmetric.
Proof.
First, from (3.3), we compute
$p^{d}Du(x)=$ $\sum_{z\neq x}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x))$ (3.6)
$=$ $\sum_{z\neq x}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}u(z)-\sum_{z\neq x}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}u(x)$
$=$ $\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
$-(\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j})u(x)$
Since $a_{0}\neq b_{0}$ correspond to $(p-2)p^{d-1}$ cases and $a_{i}=b_{i},\ i=0,1,\dots,j-1,\ a_{j}\neq b_{j}$ has $(p-1)p^{d-1-j}$ cases for $1\leq j\leq d-1$, so the coefficient of $u(x)$ is
$\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}$ (3.7)
$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d}\sum_{i=0}^{d-2}p^{i}$
$=$ $(p-1)p^{d-1}\frac{2-|x|_{p}-p^{1-m}|x|_{p}^{-1}}{p-1}+(p-2)p^{d-1}+p^{d}(p^{d-1}-1)$
$=$ $(2-|x|_{p}-p^{1-m}|x|_{p}^{-1})p^{d-1}+p^{2d-1}-2p^{d-1}$
$=$ $p^{2d-1}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1}$
which coincides with the negative of the diagonal entries of $\boldsymbol{B}$. Note that
$\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{2(d-1)}u(x)$
$=$ $\sum_{\begin{subarray}{c}a_{0}\neq b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}p^{2}u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\neq b_{d-1}\end{subarray}}p^{2(d-1)}u(z)$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-1)u(z)$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}=b_{1}\\ a_{2}\neq b_{2}\\ a_{3}\dots a_{d-1}\end{subarray}}(p^{4}-p^{2})u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2})u(z)$
$=$ $\cdots$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2(d-2)})u(z)$
$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$
so we have
$p^{d}Du(x)=$ $\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)$
$+\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$
$-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)$
then when $x\neq y$ the equation (3.2) becomes
$\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$ (3.8)
$+\rho p^{d}e^{u(x)}-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)=0$
Denote $u(x)=u_{l}(b_{0},b_{1},\dots,b_{d-1})$, by fixing $b_{0},b_{1},\dots,b_{d-2}$ and varying $b_{d-1}$, we see that $u(x)$ satisfies (LABEL:1_4). In addition, the function $f(x)=\rho p^{d}e^{x}-kx,\ x\leq d\text{ln}p$ is monotonic when $p^{2d}\rho\leq k$. Then by Lemma 3.2, if $p^{2d}\rho\leq p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-1})$ is constant for all $b_{d-1}$ except at $x=y$. Hence for $x\neq y$, we denote $u(x)$ by $u_{l}(b_{0},b_{1},\dots,b_{d-2})$.
When $v_{p}(x)=l\neq v_{p}(y)$, (LABEL:1_4) becomes
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{v_{p}(z)=l}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$
$+p(p^{2(d-1)}-p^{2(d-2)})u_{l}(b_{0},b_{1},\dots,b_{d-2})+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}$
$-(p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0$
This simplifies to
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{v_{p}(z)=l}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$ (3.9)
$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0$
Fixing $b_{0},b_{1},\dots,b_{d-3}$ and varying $b_{d-2}$, we see that $u(x)$ satisfies (LABEL:1_5). So when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-2})$ is constant for all $b_{d-2}$, denote it by $u_{l}(b_{0},b_{1},\dots,b_{d-3})$. Similarly, if we fix $b_{0},b_{1},\dots,b_{d-4}$ and when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-3})$ is constant for all $b_{d-3}$. Proceeding recursively, we obtain
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{z}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},\dots,a_{s},a_{s+1})$ (3.10)
$+\rho p^{d}e^{u_{l}(b_{0},\dots,b_{s},b_{s+1})}-(p^{d+1+s}+p^{2d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},\dots,b_{s},b_{s+1})=0$
$\forall s=0,1,\dots,d-2$. So when $\rho p^{2d}\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}$ dependent only on $b_{0},b_{1},\dots,b_{s}$, $\forall s=0,1,\dots,d-2$. In conclusion, when $v_{p}(x)=l\neq v_{p}(y)$, if $\rho p^{2d}\leq p^{d}+p^{d-1}-(p^{-l}+p^{l+1-m})p^{d-1}$, then $u_{l}$ is constant.
When $v_{p}(x)=l=v_{p}(y)$, denote $y=p^{l}(c_{0}+c_{1}p+\dots+c_{d-1}p^{d-1})$, since $u_{l}(b_{0},b_{1},\dots,b_{d-1})$ is constant for all $b_{d-1}$ except at $x=y$, fixing $b_{i}=c_{i}$, $0\leq i\leq d-2$, $b_{d-1}\neq c_{d-1}$, then $u_{l}(c_{0},\dots,c_{d-2},b_{d-1})$ is independent of $b_{d-1}$, denote it by $u_{l}(c_{0},\dots,c_{d-2})$.
If we fix $b_{i}=c_{i}$, $0\leq i\leq s$, $b_{s+1}\neq c_{s+1}$ and suppose $u_{l}(c_{0},\dots,c_{s},b_{s+1},\dots,b_{d-1})$ is independent of $b_{s+2},\dots,b_{d-1}$, denote it by $u_{l}(c_{0},\dots,c_{s},b_{s+1})$. Next, we prove that $u_{l}(c_{0},\dots,c_{s},b_{s+1})$ is independent of $b_{s+1}$: From (3.6) and (LABEL:coefficient_of_u(x)), we can get
$p^{d}Du(x)=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.11)
$-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
$+\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
we compute
$\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
$=$ $p^{2(s+1)}(p^{d-s-2}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+(p-1)p^{d-s-3}u_{l}(c_{0},\dots,c_{s+1})+\dots+(p-1)u_{l}(c_{0},\dots,c_{d-2})+u(y))$
$=$ $p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
and
$\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)=$ $\sum_{j=s+2}^{d-1}(p-1)p^{d-1-j}p^{2j}u_{l}(c_{0},\dots,c_{s},b_{s+1})$
$=$ $(p^{2d-1}-p^{d+s+1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
substituting into (3.11) yields
$p^{d}Du(x)$ (3.12)
$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)-(p^{d+s+1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
so when $p^{2d}\rho\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}$ depend only on $c_{0},c_{1},\dots,c_{s}$, $\forall s=0,1,\dots,d-2$. Then we get the result by induction. In conclusion, when $v_{p}(x)=l=v_{p}(y)$, $p^{2d}\rho\leq p^{d+1+s}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(c_{0},\dots,c_{s})$ is constant. Then when
$\rho$ $\leq p^{-2d}\text{min}_{\begin{subarray}{c}-1\leq s\leq d-2\\ 0\leq l\leq m-1\end{subarray}}\{p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}\}$
$=p^{-2d}(p^{d}+p^{d-1}-(1+p^{1-m})p^{d-1})$
$=p^{-2d}(p^{d}-p^{d-m})$
$=p^{-d}-p^{-d-m}$
$u(x)$ is radial symmetric. By Lemma 3.4 and Lemma 3.5, we can see that $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, and if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ with $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$, which means that $u(x)$ is radial symmetric. ∎
From our numerical results, we can also see that the radial symmetric solutions have monotonicity. $u(x)$ increases with the distance from the singular point $y$.
Theorem 3.7.
Assume $\rho\leq p^{-d}-p^{-d-m}$ and $v_{p}(x)=l=v_{p}(y)$, then $u_{l}(c_{0},\dots,c_{d-1})<u_{l}(c_{0},\dots,c_{d-2})<\dots<u_{l}(c_{0})<u_{l}$
Proof.
Let $v_{p}(z)=k$, by Theorem 3.6, when $\rho\leq p^{-d}-p^{-d-m}$, $u(x)$ is radial symmetric, so the equation variables can be reduced from $m(p-1)p^{d-1}$ to $p+m$. Then denote these variables by $u_{0},\dots,u_{l-1},u_{l},u_{l}(c_{0}),\dots,u_{l}(c_{0},\dots,c_{d-1}),u_{l+1},\dots,u_{m-1}$.
When $x\neq y$, equation (3.4) becomes
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}+p^{d}\rho e^{u_{l}(x)}=0$ (3.13)
When $b_{i}=c_{i}$ for $0\leq i\leq s$, then from (LABEL:Du_3), we have
$\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s})$ (3.14)
$=$ $\sum_{j=0}^{s}\sum_{a_{j}\neq c_{j}}p^{d-1-j}p^{2j}u_{l}(c_{0},\dots,c_{j-1})+p^{d+s}\sum_{a_{s+1}\neq c_{s+1}}u_{l}(c_{0},\dots,c_{s})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})$
$+p^{2(s+1)}u_{l}(c_{0},\dots,c_{d-1})-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
$=$ $\sum_{j=0}^{s}\sum_{a_{j}\neq c_{j}}p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{d+s}(p-1-\delta_{s,-1})u_{l}(c_{0},\dots,c_{s})$
$+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
$=$ $\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
$-(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
where $s=-1,0,1,\dots,d-2$.
Next, we will use backward mathematical induction to prove this theorem:
From Lemma 3.3, we know that $u(y)$ is the only smallest solution, so we have $u_{l}(c_{0},\dots,c_{d-2})>u_{l}(c_{0},\dots,c_{d-1})$.
Suppose $u_{l}(c_{0},\dots,c_{s+1})>\dots>u_{l}(c_{0},\dots,c_{d-1})$, now we want to compare $u_{l}(c_{0},\dots,c_{s+1})$ with $u_{l}(c_{0},\dots,c_{s})$. First, by (LABEL:eq2_1) and (LABEL:eq2_2), they satisfy the following equations:
$(p-1)p^{d-1}(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k})+\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.15)
$+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
$-(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s})}=0$
$(p-1)p^{d-1}(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k})+\sum_{j=0}^{s+1}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.16)
$+p^{2(s+2)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
$-(2p^{s+2}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}=0$
subtracting (LABEL:eq2_4) from (LABEL:eq2_3) gives
$0=$ $p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}-p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s})}$
$-(2p^{s+2}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})-p^{2(s+1)}(p-1)p^{d-2-s-1}u_{l}(c_{0},\dots,c_{s+1})$
$+(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
$+(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+(p-1-\delta_{s+1,0})p^{d+s}u_{l}(c_{0},\dots,c_{s})$
which is
$0=$ $p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}-p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s})}$
$-(2p^{s+2}+p^{s+1}-p^{s}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})$
$+(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
then by induction hypothesis and noting that
$(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}+1)$
$=$ $(p^{2}-1)p^{2(s+1)}((p-1)p^{d-2}\frac{p^{-s-1}-p^{-d+2}}{p-1}+1)$
$=$ $p^{d-1}(p^{s+2}-p^{s})$
we obtain
$0<$ $p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}-p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s})}$
$-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})$
$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
which is
$0<$ $\rho e^{u_{l}(c_{0},\dots,c_{s+1})}-\rho e^{u_{l}(c_{0},\dots,c_{s})}$
$-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s+1})$
$+p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
since $\rho\leq p^{-d}-p^{-d-m}$, then by the monotonicity of $f_{s}(x)=\rho e^{x}-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})x$, we have $u_{l}(c_{0},\dots,c_{s+1})<u_{l}(c_{0},\dots,c_{s})$.
Remark 3.8.
The differences $u_{l}(c_{0},\dots,c_{s})-u_{l}(c_{0},\dots,c_{s+1})$decrease with $s$ as
$f_{s}(u_{l}(c_{0},\dots,c_{s+1}))-f_{s}(u_{l}(c_{0},\dots,c_{s}))$
$=$ $(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}((u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{j}))$
$+(u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{d-1}))$
so by the monotonicity of the solution and $f_{s}(x)$, the gap between them decreases.
This theorem provides a sharper estimate of the upper bound for the solution. We now refine Lemma 3.2:
Lemma 3.9.
$u(i)<\begin{cases}\text{ln}\frac{p}{p-2},\ &p>2\\ 4,&p=2\end{cases}$.
Proof.
From Lemma 3.2, we know that $\sum_{i}e^{u_{i}}=p^{d}$. Denoting $M$ to be the largest solution. Then from Theorem 3.7, we can see there at least $(p-2)p^{d-1}$ such $M$ when $p>2$, so we get
$(p-2)p^{d-1}e^{M}<p^{d}$
thus $M<\text{ln}\frac{p}{p-2}$.
When $p=2$, there are at least $2^{d-2}$ such $M$, so $M<4$. ∎
Using this improved bound, we refine Theorem (3.6) and Theorem (3.7):
Theorem 3.10.
When $\rho\leq\begin{cases}\frac{p-2}{p}(1-p^{-m}),\ &p>2\\ \frac{1-2^{-m}}{4},&p=2\end{cases}$, then the solution is radial symmetric and monotone.
3.2 Uniqueness of the solution
In this part, we discuss the uniqueness of the solution. Numerical experiments indicates that for large $\rho$, solutions may not unique, and we can get some solutions that break the properties of radial symmetric. Then by Lemma 3.4 and Lemma 3.5, we can construct many non-radial symmetric solutions. A natural question is whether uniqueness holds under the conditions of $\rho$ guaranteeing radial symmetric of the solution. The answer is affirmative with a slighter stricter condition on $\rho$.
Uniqueness is straightforward on the finite quotient. If $\boldsymbol{u}$ and $\boldsymbol{v}$ both solve (3.4). Then
$\boldsymbol{A}(\boldsymbol{u}-\boldsymbol{v})+\rho(e^{\boldsymbol{u}}-e^{\boldsymbol{v}})=\boldsymbol{0}$
By the mean value theorem applied for each component, there exists $\boldsymbol{\xi}$ such that
$(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})(\boldsymbol{u}-\boldsymbol{v})=\boldsymbol{0}$
Lemma 3.9 gives $e^{\xi_{i}}\leq\frac{p}{p-2}$, which is independent of $\rho$. Since $\boldsymbol{A}$ is semi-negative definite, if $\rho\frac{p}{p-2}$ does not exceed the first non-zero eigenvalue of $\boldsymbol{A}$, then $(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})$ is invertible, then we get the uniqueness.
So the next step is to study the eigenvalue of $\boldsymbol{A}$, i.e. the spectrum of $D$ on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$. As we can pull back the functions from $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ to $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$, so the eigenvalue of $\boldsymbol{A}$ is just the subset of the spectrum of $D$ by taking the first $d$-th eigenvalues from each component. Alternatively, we can compute the eigenvalues of $\boldsymbol{A}$ via linear algebra as follows:
Theorem 3.11.
Eigenvalues and the eigenvectors of $\boldsymbol{A}$.
Proof.
As $\boldsymbol{A}=p^{-d}\boldsymbol{B}$, we first consider $\boldsymbol{B}$. Denoting $n=(p-1)p^{d-1}$. Then $\boldsymbol{B}\in\mathbb{R}^{mn\times mn}$. From previous observation, $\boldsymbol{B}$ is a block matrix and each block is a circulant matrix. Furthermore, the off-diagonal part is a matrix with the same element. Denoting $\boldsymbol{P}\in\mathbb{R}^{m\times m}$ with $(\boldsymbol{P})_{ij}=p^{-|i-j|}$. And $\boldsymbol{T}\in\mathbb{R}^{n\times n}$ be the circulant part on diagonal with $(\boldsymbol{T})_{ij}=(\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}-1)(1-\delta_{zx})$ (i.e. $\boldsymbol{T}$ is zero on diagonal and $\boldsymbol{T}$ is independent of $v_{p}(z)=v_{p}(x)$). Then we can write $\boldsymbol{B}$ as
$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-\text{diag}((\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn})$
First, we calculate
$(\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn}=$ $(\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}$
$=$ $\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}\boldsymbol{1}_{n}+\boldsymbol{I}_{m}\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n}$
$=$ $n\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}+\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n}$
let $t_{0}$ be the row sum of $\boldsymbol{T}$. Then
$\text{diag}((\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn})=$ $\text{diag}(n\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n})+\text{diag}(\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n})$
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n})+\text{diag}(\boldsymbol{1}_{m}\otimes t_{0}\boldsymbol{1}_{n})$
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn}$
hence
$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn})$
Let $\boldsymbol{F}$ be the DFT matrix, as any circulant matrix can be diagonalized by a DFT matrix, denoting the eigenvalue of $\boldsymbol{T}$ is $t_{0},t_{1}\dots,t_{n-1}$, which is real and $t_{0}$ is the largest as $\boldsymbol{T}$ is a symmetric circulant matrix and the elements in $\boldsymbol{T}$ are nonnegative. Then
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=$ $\boldsymbol{I}_{m}\boldsymbol{P}\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}\boldsymbol{F}$
$=$ $\boldsymbol{P}\otimes\begin{pmatrix}n&\\ &\boldsymbol{O}_{n-1}\end{pmatrix}$
$=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &\boldsymbol{O}_{mn-m}\end{pmatrix}$
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}\boldsymbol{T}\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda}$
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{F}^{-1}\boldsymbol{I}_{n}\boldsymbol{F}$
$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}$
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(t_{0}\boldsymbol{I}_{mn})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=t_{0}\boldsymbol{I}_{mn}$
Thus
$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{B})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &0\end{pmatrix}+\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}-t_{0}\boldsymbol{I}_{mn}$
$=$ $\begin{pmatrix}n\boldsymbol{P}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})&\\ &(t_{i}-t_{0})\boldsymbol{I}_{m}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\end{pmatrix}$
Denoting $s_{j}$ the row sum of $\boldsymbol{P}$, we have
$s_{j}=$ $\sum_{i=0}^{m-1}p^{-|i-j|}$
$=$ $\sum_{i=0}^{j-1}p^{-(j-i)}+\sum_{i=j}^{m-1}p^{-(i-j)}$
$=$ $\frac{1+p-p^{-j}-p^{j+1-m}}{p-1}$
So the eigenvalue of $\boldsymbol{B}$ is $n\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))=(p-1)p^{d-1}\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$ and $t_{i}-t_{0}-ns_{j}=t_{i}-t_{0}-p^{d-1}(1+p-p^{-j}-p^{j+1-m})$, where $i=1,\dots,n-1;\ j=0,\dots,m-1$. Hence, the eigenvalue of $\boldsymbol{B}$ is $(1-p^{-1})\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$ and $p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}$. Note that $\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ is symmetric, then there exists an orthogonal matrix $\boldsymbol{Q}\in\mathbb{R}^{m\times m}$ to diagonalize it. So the corresponding eigenvector is $\boldsymbol{I}_{m}\otimes\boldsymbol{F}\begin{pmatrix}\boldsymbol{Q}&\\ &\boldsymbol{I}_{m(p-1)p^{d-1}-m}\end{pmatrix}$. Moreover, the spectrum of $\boldsymbol{A}$ must be compatible with the spectrum of $D$. So by Theorem 2.4 we have $p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}=-p^{i-1}-p^{i-2}+\frac{p^{-j}+p^{j+1-m}}{p}$. In summary, we get
$(\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1})\begin{pmatrix}\boldsymbol{Q}^{-1}&\\ &\boldsymbol{I}_{m(p-1)p^{d-1}-m}\end{pmatrix}(\boldsymbol{A})(\boldsymbol{I}_{m}\otimes\boldsymbol{F})\begin{pmatrix}\boldsymbol{Q}&\\ &\boldsymbol{I}_{m(p-1)p^{d-1}-m}\end{pmatrix}$
$=$ $\begin{pmatrix}(1-p^{-1})\lambda_{k}&\\ &-p^{i-1}-p^{i-2}+\frac{p^{-j}+p^{j+1-m}}{p}\end{pmatrix}$
where $\lambda_{k}$ are the eigenvalues of $\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ for $k=0,\dots,m-1$, and $i=1,\dots,d-1$, $j=0,\dots,m-1$. ∎
Lemma 3.12.
The smallest non-zero eigenvalue of $\boldsymbol{A}$ is $\begin{cases}(1-p^{-1}),&m=1\\ -(1-p^{-1})\lambda_{1},&m>1\end{cases}$
Proof.
Denoting $\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})=-\boldsymbol{L}$, then the eigenvalues of $\boldsymbol{L}$ is $\mu_{i}=-\lambda_{i}$ with $0=\mu_{0}<\mu_{1}<\dots<\mu_{m-1}$. As $\mu_{0}=0$ with eigenvector $\boldsymbol{1}_{m}$, so by Courant–Fischer min-max theorem, we have
$\mu_{1}=\text{min}_{\begin{subarray}{c}\boldsymbol{x}\neq\boldsymbol{0}_{m}\\ \boldsymbol{x}\bot\boldsymbol{1}_{m}\end{subarray}}\frac{\boldsymbol{x}^{\top}\boldsymbol{L}\boldsymbol{x}}{\boldsymbol{x}^{\top}\boldsymbol{x}}$
taking $\boldsymbol{x}=\boldsymbol{e}_{0}-\boldsymbol{e}_{m-1}$, we have
$\frac{\boldsymbol{x}^{\top}\boldsymbol{L}\boldsymbol{x}}{\boldsymbol{x}^{\top}\boldsymbol{x}}=$ $\frac{1}{2}(L_{0,0}+L_{m-1,m-1}-2L_{0,m-1})$
$=$ $\frac{1}{2}(2\frac{1-p^{1-m}}{p-1}+2p^{1-m})$
$=$ $\frac{1+(p-2)p^{1-m}}{p-1}$
$\leq$ $1$
with equality if and only if $p=2$.
Therefore, the smallest nonzero eigenvalue of $\boldsymbol{A}$ is
$\text{min}_{\begin{subarray}{c}1\leq k\leq m-1\\ 1\leq i\leq d-1\\ 0\leq j\leq m-1\end{subarray}}\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\}$
$=$ $\text{min}\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
$=$ $\begin{cases}(1-p^{-1}),&m=1\\ -(1-p^{-1})\lambda_{1},&m>1\end{cases}$
Theorem 3.13.
When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then the solution of (3.2) is unique.
Proof.
By Lemma 3.9, when $p>2$, we have
$\rho\frac{p}{p-2}\leq(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$
thus $\rho\leq(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$.
When $p=2$,we have
$4\rho\leq\frac{1}{2}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$
thus $\rho\leq\frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$. ∎
4 Mean field equation on the Tate curve
Having studied the solution on finite quotient, we now consider the solution on the Tate curve. Since the finite quotient converges to the Tate curve when $d\rightarrow\infty$, we investigate whether the solution converge accordingly. The answer is affirmative. Moreover, the limiting solution retains properties like uniqueness and monotonicity.
4.1 Convergence to the Tate curve
In this part, we will show that the solution on the finite quotient will converge to the solution on the Tate curve, thus we can get the existence of the solution. From the numerical experiments, we can indeed observe that the solution on the finite quotient is convergent. Next, we will prove it rigorously.
Theorem 4.1.
When $\rho\leq\begin{cases}\frac{p-2}{p}(1-p^{-m}),\ &p>2\\ \frac{1-2^{-m}}{4},&p=2\end{cases}$, then there exists a solution $u$ satisfying (3.1).
Proof.
For convenience, we denote $X_{m}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with the Haar measure $d\mu=\frac{dx}{|x|_{p}}$ and $X_{m,d}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ with the Haar measure $d\mu_{d}=p^{-d}d\delta$. Clearly, $\mu_{d}\overset{*}{\rightharpoonup}\mu$ as $d\rightarrow\infty$. Let $D_{d}$ denote the operator on the finite quotient. Consider
$D_{d}u_{d}+\rho e^{u_{d}}=\rho p^{d}e_{y}\ \text{in}\ X_{m,d}$ (4.1)
with
$\int_{X_{m,d}}e^{u_{d}}d\mu_{d}=1$ (4.2)
Since $D_{d}u_{d}=\rho p^{d}e_{y}-\rho e^{u_{d}}$, we have $|D_{d}u_{d}|\leq\rho p^{d}e_{y}+\rho e^{u_{d}}$, then
$\int_{X_{m,d}}|D_{d}u_{d}|d\mu_{d}\leq 2\rho$ (4.3)
Define measures $\nu_{d}=e^{u_{d}}\mu_{d}$ and $\omega_{d}=D_{d}u_{d}\mu_{d}$ on $X_{m,d}$. As $X_{m,d}\hookrightarrow X_{m}$, so they can be seen as the restriction of measure $\widetilde{\nu_{d}}$ and $\widetilde{\omega_{d}}$ on $X_{m}$, i.e. $\widetilde{\nu_{d}}|_{X_{m,d}}=\nu_{d}$ and $\widetilde{\omega_{d}}|_{X_{m,d}}=\omega_{d}$.
Since $X_{m}$ is compact, we have $C(X_{m})^{*}=BV(X_{m})$, so by the Banach-Alaoglu theorem, $BV(X_{m})$ is weak-* compact. Then from (4.2) and (4.3), we have $\widetilde{\nu_{d}}||_{BV}=1$ and $\widetilde{\omega_{d}}||_{BV}\leq 2\rho$, so there exists $\nu$, $\omega\in BV(X_{m})$ and a subsequence $d_{k}$ such that
$\widetilde{\nu_{d_{k}}}\overset{*}{\rightharpoonup}\nu,\ \widetilde{\omega_{d_{k}}}\overset{*}{\rightharpoonup}\omega$
as $\nu_{d}<<\mu_{d}$ and $\frac{d\nu_{d}}{d\mu_{d}}=e^{u_{d}}\leq C_{p}$ uniformly in $d$ by Lemma 3.9, we have $\nu<<\mu$, denoting $\nu=e^{\widetilde{u}}\mu$ for some function $\widetilde{u}$.
Take any locally constant function $\phi$ on $X_{m}$, so when $d$ is large enough, $\phi$ is well-defined on $X_{m,d}$ as $X_{m}$ is compact. Multiply (4.1) by $\phi$ and integrate over $X_{m,d}$, we have
$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}+\rho\int_{X_{m,d}}e^{u_{d}}\phi d\mu_{d}=\rho\phi(y)$
by taking a limit of the subsequence $d_{k}$, we get
$\int_{X_{m}}\phi d\omega+\rho\int_{X_{m}}\phi d\nu=\rho\phi(y)$ (4.4)
Decompose $u_{d}=v_{d}+\rho G_{d}$, where $G_{d}$ is the Green’s function on finite quotient, then $v_{d}$ satisfies
$D_{d}v_{d}=-\rho e^{u_{d}}+\frac{\rho}{V}=-\rho e^{\rho G_{d}}e^{v_{d}}+\frac{\rho}{V}$
so
$v_{d}(x)=\int_{X_{m,d}}G_{d}(x,y)(-\rho e^{u_{d}(y)}+\frac{\rho}{V})d\mu_{d}(y)$
Since $G(x,y)\sim\frac{p(p-1)}{p+1}\text{log}_{p}d(x,y)$ as $d(x,y)\rightarrow 0$ [14], and $X_{m}$ is compact, then
[…]
and we have $\int_{B_{\epsilon}(x)}|G(x,y)|d\mu(y)\leq C\epsilon|\text{ln}\epsilon|$, so $G\in L^{1}(X_{m})$. and we notice that $\rho e^{u_{d}(y)}+\frac{\rho}{V}$ is bounded, we have
[…]
so $v_{d}\in L^{\infty}(X_{m,d})$.
Using the self-adjointness of $D$, we have
$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}=$ $\int_{X_{m,d}}(D_{d}\phi)u_{d}d\mu_{d}$
[…]
$=$ $\mathrm{I}+\mathrm{II}$
where
$|\mathrm{I}|=$ $\left|\int_{B_{\epsilon}(y)}(D_{d}\phi)u_{d}d\mu_{d}\right|$
$\leq$ $\rho\int_{B_{\epsilon}(y)}|D_{d}\phi(x)||G_{d}(x,y)|d\mu_{d}(x)+\int_{B_{\epsilon}(y)}|D_{d}\phi||v_{d}|d\mu_{d}$
$\leq$ $C_{1}\epsilon|\text{ln}\epsilon|+C_{2}\epsilon$
and we have
$\mathrm{II}=\int_{X_{m,d}\backslash B_{\epsilon}(y)}(D_{d}\phi)u_{d}d\mu_{d}=\int_{X_{m,d}\backslash B_{\epsilon}(y)}(D\phi)u_{d}d\mu_{d}$
for sufficiently large $d$. And in $X_{m,d}\backslash B_{\epsilon}(y)$, $|u_{d}|\leq C_{\epsilon}$, so there exists a subsequence $u_{d_{k}}\rightarrow u$ such that
$\mathrm{II}\rightarrow\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\mu$
letting $\epsilon\rightarrow 0$ we obtain
$\int_{X_{m,d_{k}}}(D_{d_{k}}u_{d_{k}})\phi d\mu_{d_{k}}\rightarrow\int_{X_{m}}(D\phi)ud\mu=\int_{X_{m}}(Du)\phi d\mu$
as $k\rightarrow\infty$. Hence $\omega=Du\mu$ in the sense of distributions.
What’s more,
$\int_{X_{m}}\phi d\nu=\lim_{k\to\infty}\int_{X_{m,d_{k}}}\phi e^{u_{d_{k}}}d\mu_{d_{k}}=\int_{X_{m}}\phi e^{u}d\mu$
and by definition,
$\int_{X_{m}}\phi d\nu=\int_{X_{m}}\phi e^{\widetilde{u}}d\mu$
so we have $\widetilde{u}=u$ for a.e. $x$. Then by (4.4) we get
$\int_{X_{m}}(Du)\phi d\mu+\rho\int_{X_{m}}e^{u}\phi d\mu=\rho\phi(y)$
[…]
Since the solution we found converges from the solution on finite quotient, it inherits the properties of the solution on finite quotient. So the solution $u$ has those radial symmetric properties, what’s more, $u$ is monotone with respect to $d(x,y)$ at the same connected component and $u\leq C_{p}$.
When $\rho$ breaks the bound of Theorem 4.1, then there may be non-radial symmetric solutions, so that $u$ will not have the universal upper bound, this may prevent the convergence of solutions on the finite quotient, and hence the existence cannot be guaranteed.
4.2 Uniqueness of the solution
Similarly from the argument of the solution in finite quotient, we now address uniqueness on the Tate curve under the same constraints of $\rho$. Although the solution on finite quotient is unique when $\rho$ is not too large, we can’t ensure the solution on the Tate curve is unique as there may has another solution on the Tate curve but vanishes on finite quotient.
Actually, the solution is indeed unique under the same constraints of $\rho$. We will prove it by a standard method [18]. First, we show that the solution is unique when $\rho$ is very small, next we prove the linearized solution is non-degenerate under the same constraints of $\rho$, then we can get the uniqueness.
Lemma 4.2.
There exists $\varepsilon>0$, such that when $\rho\leq\varepsilon$, then (3.1) has a unique solution.
Proof.
Let $u_{n}$, $v_{n}$ be distinct solutions of (3.1) with $\rho_{n}\rightarrow 0$. Since $\int_{\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}}e^{u}\frac{dx}{|x|_{p}}=1$, the difference $u_{n}-v_{n}$ must change sign. Define
$\phi_{n}=\frac{u_{n}-v_{n}}{\|u_{n}-v_{n}\|_{L^{\infty}}}$
Then $\phi_{n}$ satisfies
$D\phi_{n}+V_{n}(x)\phi_{n}=0$
where $V_{n}(x)=\rho_{n}\frac{e^{u_{n}}-e^{v_{n}}}{u_{n}-v_{n}}$. Since $u(x)$ has a universal upper bound by Lemma 3.9, $V_{n}(x)$ converges to 0. And it is easy to see that $\phi_{n}$ uniformly converges to the eigenfunction of $D$ with the eigenvalue 0, which is the harmonic function on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$. As $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ is compact without boundary, then the harmonic function on it must be a constant function, which contradicts the fact that $\phi_{n}$ must change the sign. ∎
Next, we consider the linearized equation of (3.1) as
$Df+\rho e^{u}f=0$ (4.5)
If (4.5) is nondegenerate for $\rho\leq C$, then Lemma 4.2 implies the uniqueness of the solution for such $\rho$. To address this question, we give the following Lemma.
Lemma 4.3.
Let $L$ be an operator with eigenvalues $\lambda_{1}\leq\lambda_{2}\leq\cdots$, and $\exists\ k$ s.t. $\lambda_{k}<V(x)<\lambda_{k+1}$. Then the operator LV(x)L-V(x) is nondegenerate.
Proof.
Suppose that the eigenvalue of LV(x)L-V(x) is $\mu_{1}\leq\mu_{2}\leq\cdots$. Then we have operator inequalities
$L-\lambda_{k+1}<L-V(x)<L-\lambda_{k}$
then we have
$\lambda_{i}-\lambda_{k+1}<\mu_{i}<\lambda_{i}-\lambda_{k}$
for all $i$. Taking $i=k$ and $i=k+1$, we get
$\mu_{k}<0<\mu_{k+1}$
which means 0 is not the eigenvalue of LV(x)L-V(x). ∎
By Lemma 3.9, we can see that $0<\rho e^{u}<\rho\frac{p}{p-2}$. And from Lemma 4.3, we can see that if $\rho\frac{p}{p-2}$ is smaller than the first nonzero eigenvalue of D-D, then we can get the uniqueness, similarly, the smallest nonzero eigenvalue of D-D is
$\text{min}_{\begin{subarray}{c}1\leq k\leq m-1\\ i\geq 1\\ 0\leq j\leq m-1\end{subarray}}\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\}$ (4.6)
$=$ $\text{min}\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
[…]
then we have
Theorem 4.4.
When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then the solution of (3.1) is unique.
Acknowledgement
I am deeply indebted to An Huang for his patient guidance, insightful discussions, and unwavering support throughout this research. I extend my sincere gratitude to my supervisor Bobo Hua for his valuable suggestions and help.

### 08 Apr 2026 v2

@@
-Mean field equation and Green’s function on the Tate curve
+The Mean field equation on the Tate curve
 Abstract
 In this paper, we study the spectrum of the Laplacian on the Tate curve and construct the associated Green’s function as a finite sum, which can be viewed as the non-Archimedean counterpart of the Green’s function on the flat torus in the Archimedean case.
-And we establish the existence and uniqueness results of the mean field equation on this space.
-To address this problem, we begin by presenting the solution structure and proving its uniqueness under specific conditions on the finite quotient.
-We then establish convergence of the solutions to obtain existence results.
+Moreover, we establish existence and uniqueness results of the mean field equation on this space.
+To address the problem, we first prove the structure of solutions on finite quotients, and prove the existence on the Tate curve by the convergence of such solutions.
+We also prove the uniqueness of the solutions for some parameter region.
 Notably, the well-posedness of the solution resembles that in the Archimedean case.
 1 Introduction
-The mean field equation
+The mean field equation,
 $\Delta u+\rho e^{u}=\rho\delta_{0},\ \rho>0$ (1.1)
-arises from statistical mechanics and resembles the Liouville equation, the latter stems from the prescribed Gaussian curvature problem but formulated without a point source term on the right-hand side.
-A more general formulation is given by the Kazdan-Warner equation.
+rooted in statistical mechanics, is formally analogous to the Liouville equation; however, the latter originates from the prescribed Gaussian curvature problem and is typically formulated without a point source term on the right-hand side.
+A more general formulation is given by the Kazdan-Warner equation [17].
 When $\rho=4\pi$, (1.1) is related to the Chern-Simons model [4].
-On special domains such as $\Omega\subseteq\mathbb{R}^{n}$ or compact manifolds, the existence and uniqueness of the solution has been studied in [5, 21, 22, 6, 7, 17].
-The existence of the solution on graphs has also been investigated.
-Grigor’yan, Lin, and Yang [9] provided necessary conditions for the existence of solutions on finite graphs, while Huang-Lin-Yau [13] established the existence of solutions on finite weighted graphs.
-Additionally, Ge and Jiang [8] further proved the existence of solutions on infinite graphs.
-Ling and Wang [20, 19] studied the mean field equation on the flat torus $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ for $\tau\in\mathbb{C}\backslash\mathbb{R}$.
-When $\rho\neq 8m\pi$ with $m\in\mathbb{Z}$, there always exist a solution as the Leray-Schauder degree is nonzero, independent of the torus shape.
-When $\rho=8\pi$, they demonstrated that the existence and uniqueness of the solution is equivalent to the number of critical points of the Green’s function
-$\Delta G(z,w)=\delta_{w}(z)-\frac{1}{V}$ (1.2)
-where $V$ denotes the volume of the torus. Specifically, if the mean field equation admits a unique solution, then the Green’s function possesses five critical points, otherwise it has only three.
+On special domains such as $\Omega\subseteq\mathbb{R}^{n}$ or compact manifolds, the existence and uniqueness of the solution have been studied in [5, 23, 24, 6, 7, 18].
+The existence of solutions to these type of equations on graphs has been investigated.
+Grigor’yan, Lin, and Yang [9] provided various conditions for the existence of the Kazdan-Warner equation on finite graphs, Ge and Jiang [8] further proved the existence of solutions on infinite graphs.
+Huang-Lin-Yau [13] established the existence of the mean field equation on finite weighted graphs via the calculus of variations.
+Further Lin and Yang introduced a mean-field type heat flow on finite graphs [25].
+In the same year, Liu studied the mean-field equation with a positive prescribed function by using the method of Brouwer degree [27].
+Later, Sun and Wang promoted their results to the sign-changing prescribed function [28].
+After that, Liu and Zhang further developed the heat flow method and resolved the convergence problem for the mean-field equation with a sign-changing prescribed function [26].
+In recent years, Li and Zhang systematically investigated the mean-field equation with a linear term by means of variational methods [19].
+And Liu extended the results to nonlocal operators via considering the fractional Laplacian $(-\Delta)^{s}$, this work unifies and generalizes the results of [25, 28, 26] for the case $s=1$.
+Lin and Wang [22, 21] studied the mean field equation on the flat torus $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)$ for $\tau\in\mathbb{C}\backslash\mathbb{R}$.
+When $\rho\neq 8m\pi$ for $m\in\mathbb{Z}$, there always exists a solution as the Leray-Schauder degree is nonzero, which is independent of the torus shape $\tau$.
+When $\rho=8\pi$, they demonstrated that the existence and uniqueness of the solution depend on the number of critical points of the Green’s function
+$\Delta G(z,w)=\delta_{w}(z)-\frac{1}{V},$ (1.2)
+where $V$ denotes the volume of the torus.
+In particular, the mean field equation admits a unique solution if and only if the Green’s function possesses five critical points (otherwise it has only three).
 Since these Liouville-type equations have numerous applications in quantum field theory and string theory [29, 10, 1], and the Green’s function underlies the computation of Veneziano amplitudes.
 It is natural to investigate a $p$-adic analogue of this problem.
-Analyses of $p$-adic pseudo-differential operators can be found in [24].
-In the genus zero case, Huang-Stoica-Yau-Zhong [15, 16] defined the $p$-adic version of the flat Laplacian via regularizing Vladimirov derivation as
+Analysis of $p$-adic pseudo-differential operators can be found in [30].
+In the genus-zero case, Huang-Stoica-Yau-Zhong [15, 16] defined the $p$-adic version of the flat Laplacian via regularizing Vladimirov derivation as
 $D^{s}u(x)=\int_{\mathbb{Q}_{p}}\frac{u(z)-u(x)}{|z-x|_{p}^{1+s}}dz$
-and constructed the associated Green’s function, where $dz$ is the Haar measure on $\mathbb{Q}_{p}$. Besides, in the upper-half plane, Hassan [11] constructed the higher Green’s function.
-For genus one, Huang-Rohrlich-Sun-Whyman [14] defined the $p$-adic version of the flat Laplacian on the Tate curve as
-$Du(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}dz$
+and constructed the associated Green’s function, where $dz$ is the Haar measure on the $p$-adic field $\mathbb{Q}_{p}$.
+Besides, in the upper-half plane, Hassan [11] constructed the higher Green’s function.
+For the genus-one case, Huang-Rohrlich-Sun-Whyman [14] defined the $p$-adic version of the flat Laplacian on the Tate curve as
+$Du(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}dz,$
 where $|q|_{p}<1$.
-It is easy to check that the operator $D$ is self-adjoint, negative semi-definite and preserves locally constant functions under the inner product induced by the Haar measure $\frac{dx}{|x|_{p}}$ on $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$.
-They proved the Green’s function only depends on $d(x,y)=\frac{|x-y|_{p}}{\text{max}\{|x|_{p},|y|_{p}\}}$ and provided a series expression for the Green’s function by correcting the logarithmic singularity on the diagonal.
+One easily checks that the operator $D$ is self-adjoint and negative semi-definite under the inner product induced by the Haar measure $\frac{dx}{|x|_{p}}$ on $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$, which preserves locally constant functions.
+They proved the Green’s function only depends on $d(x,y)=\frac{|x-y|_{p}}{\max\{|x|_{p},|y|_{p}\}}$ and provided a series expression for the Green’s function by correcting the logarithmic singularity on the diagonal.
 Remark 1.1.
@@
 In the Archimedean case, we consider the exponential map: $x\mapsto e^{2\pi ix}$, yielding $\mathbb{C}/(\mathbb{Z}+\mathbb{Z}\tau)\rightarrow\mathbb{C}^{\times}/(e^{2\pi i\tau})^{\mathbb{Z}}$.
-So the Tate curve $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ with $|q|_{p}<1$ can be viewed as the non-Archimedean analogy of the flat torus.
+Based on that, the Tate curve $\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ with $|q|_{p}<1$ can be viewed as the non-Archimedean analogy of the flat torus.
 Remark 1.2.
-Since multiplication by a unit in $\mathbb{Z}_{p}^{*}$ leaves $D$ invariant, we may restrict to the case $q=p^{m}$, then we choose the fundamental domain as $\cup_{k=0}^{m-1}p^{l}\mathbb{Z}_{p}^{*}$.
+Since multiplication by a unit in $\mathbb{Z}_{p}^{*}$ leaves $D$ invariant, we may restrict to the case $q=p^{m}$, and choose the fundamental domain as $\bigcup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}$.
 One can verify that the definition is independent of the choice of the fundamental domain.
@@
 In this paper, we focus on the mean field equation and the Green’s function on the Tate curve as
-$DG(x,y)=\delta_{y}(x)-\frac{1}{V}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ (1.3)
-$Du+\rho e^{u}=\rho\delta_{y}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}$ (1.4)
-A standard approach to constructing the Green’s function employs the spectrum of the operator $D$ and its eigenfunctions.
+$DG(x,y)=\delta_{y}(x)-\frac{1}{V}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}},$ (1.3)
+$Du+\rho e^{u}=\rho\delta_{y}\ \text{in}\ \mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}.$ (1.4)
+A standard approach to construct the Green’s function employs the spectrum of the operator $D$ and its eigenfunctions.
 When $m=1$, the fundamental domain is a locally compact group, allowing us to obtain its spectrum and eigenfunctions via characters, some of these works have been established in [12].
@@
 Using orthogonality of the characters, we express the Green’s function as a finite sum, the resulting formula differs form that in [14].
-Let $\boldsymbol{L}=\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ and $\boldsymbol{R}=-(1-p^{-1})\boldsymbol{L}$, where $(\boldsymbol{P})_{ij}=p^{-|i-j|}$, so we have orthogonal matrix $\boldsymbol{Q}$ to diagonalize $\boldsymbol{R}$.
-And we define $\phi_{h,c,n}(x)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\{\frac{c\text{ln}\frac{x}{a_{0}}}{p^{n}}\}_{p}}$, where $x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, $0\leq h\leq p-2$ and $c\in\mathbb{Z}_{p}^{*}$.
-Then our first main result is:
+Letting $\boldsymbol{R}=(1-p^{-1})(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$, where $(\boldsymbol{P})_{ij}=p^{-|i-j|}$ and $\text{diag}(\boldsymbol{v})$ denotes the diagonal matrix with the vector $\boldsymbol{v}$ as its diagonal entries.
+We have orthogonal matrix $\boldsymbol{Q}$ to diagonalize $\boldsymbol{R}$, and we define
+$\phi_{h,c,n}(x)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\frac{c\ln\left(\frac{x}{a_{0}}\right)}{p^{n}}\right\}_{p}},$
+where $x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, $0\leq h\leq p-2$ and $c\in\mathbb{Z}_{p}^{*}$ with $cp^{n-1}\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$. Then our first main result is as follows:
 Theorem 1.3 (The spectrum of $D$ and the corresponding Green’s function).
@@
 when $v_{p}(x)=v_{p}(y)=l$ and $v_{p}(d(x,y))=M$.
-Inspired by the existence results for the Green’s function [14], we first consider the mean field equation on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$.
-Since there are finitely many points in the finite quotient, we can solve the equation by numerical experiments.
-Based on observations from numerical results, we investigate a universal upper bound of the solution and establish uniqueness when the parameter $\rho\leq C(p,m)$.
-We then study the convergence of the solutions to obtain existence on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$.
-Similar to the arguments on the finite quotient, the uniqueness can be preserved, yielding a unique solution.
+Inspired by the existence results for the Green’s function [14], we consider the mean field equation on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ first.
+The operator $D$ on this finite quotient simplifies to a graph Laplacian, reducing the problem to an equation on a finite weighted graph.
+Motivated by these, we prove a universal upper bound for the solution and establish its uniqueness under the condition of the parameter $\rho$.
+We then study the convergence of the solutions to obtain the existence result on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$.
+Similar to the arguments on the finite quotient, the uniqueness can be also derived.
 The optimal upper bound of $\rho$ can be viewed as analogous to the $8\pi$ threshold in Archimedean case.
-Then our second main result is:
+Then our second main result is as follows:
 Theorem 1.4.
-When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then there exists a unique solution of the mean field equation (1.4).
-Where $\lambda_{1}$ is the smallest nonzero eigenvalue of $\boldsymbol{L}$.
-Our ultimate goal is to connect the Green’s function with the mean field equation to explore the critical point properties of the Green’s function.
-Since the normal derivative on the Green’s function is trivial due to its local constancy, our first step will be to define a suitable derivative on the Green’s function in future work.
+When $\rho\in(0,\rho^{*})$, then there exists a unique solution of the mean field equation (1.4), where $\rho^{*}=\theta_{p,m}e^{-M}$ with
+[…]
+and $\theta_{p,m}$ is the smallest nonzero eigenvalue of D-D. Furthermore, the solution is radially symmetric and strictly monotone.
+Our goal is to connect the Green’s function with the mean field equation to explore the critical point properties of the Green’s function in the future works.
+Since the results in Archimedean case were obtained indirectly through the mean-field equation, our results will provide some preliminary steps toward this problem.
 1.1 Outline of the paper
@@
 In section 3, we study the mean field equation on finite quotients of the Tate curve.
-We derive a uniform positive upper bound $C_{p}$(depending only on $p$) for the solution and demonstrate its central symmetry, cyclic symmetry, and monotonicity.
-Furthermore, we prove that the solution is radial symmetric and dependents only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$ when $\rho\leq C_{p}(1-p^{-m})$, which is similar to the property of the Green’s function.
-Lastly we establish uniqueness for $\rho\leq C_{p}\mu_{1}$, where $\mu_{1}$ is the smallest nonzero eigenvalue of D-D.
+We derive a uniform positive upper bound $M$ (depending only on $p$ and $m$) for the solution and demonstrate its central symmetry, cyclic symmetry, and monotonicity.
+Furthermore, we prove that the solution is radially symmetric and dependents only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$ when $\rho\leq(1-p^{-m})e^{-M}$, which is similar to the property of the Green’s function.
+Lastly we establish uniqueness for $\rho<\rho^{*}$.
 In section 4, we show that the solutions on the finite quotient converge to a solution on the Tate curve while preserving their structural properties, thereby establishing existence, and we prove that under the same condition the solution is unique.
@@
 This section focuses on the spectrum of $D$ and its eigenfunctions.
-First, we consider the case $m=1$, then the fundamental domain is $\mathbb{Z}_{p}^{*}$, which is a compact multiplicative group.
-We analyze $D$ via characters of $\mathbb{Z}_{p}^{*}$.
-After that, we construct the Green’s function with a finite sum.
-Lastly, we generalize it to the general $m$ case.
+First, we consider the case $m=1$, where the fundamental domain is $\mathbb{Z}_{p}^{*}$, which is a compact multiplicative group.
+We analyse $D$ via characters of $\mathbb{Z}_{p}^{*}$ and construct the Green’s function with a finite sum.
+Then we generalize it to the general $m$ case through zero extension.
 2.1 Multiplicative Characters of $\mathbb{Z}_{p}^{*}$
 Since $\mathbb{Z}_{p}^{*}\cong(\mathbb{Z}/p\mathbb{Z})^{\times}\times(1+p\mathbb{Z}_{p})$, where the first factor is cyclic and the second is torsion free, it suffices to determine the characters of both components.
-Let $g$ be a primitive root of $(\mathbb{Z}/p\mathbb{Z})^{\times}$, as the dual group of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ is isomorphic to itself, so for $m,h\in\{0,1,\dots,p-2\}$, the characters for $(\mathbb{Z}/p\mathbb{Z})^{\times}$ are given by
-$\chi_{h}(g^{m})=e^{\frac{2\pi ih}{p-1}m}$
+Let $g$ be a primitive root of $(\mathbb{Z}/p\mathbb{Z})^{\times}$, as the dual group of $(\mathbb{Z}/p\mathbb{Z})^{\times}$ is isomorphic to itself, so for $w,h\in\{0,1,\dots,p-2\}$, the characters for $(\mathbb{Z}/p\mathbb{Z})^{\times}$ are given by
+$\chi_{h}(g^{w})=e^{\frac{2\pi ih}{p-1}w}.$
 As the dual group of $\mathbb{Z}_{p}$ is isomorphic to $\mathbb{Q}_{p}/\mathbb{Z}_{p}$, so the characters on $\mathbb{Z}_{p}$ are $\chi_{t}(x)=e^{2\pi i\{tx\}_{p}}$, where $x\in\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$.
-And we notice that there is a continuous isomorphism $\mathbb{Z}_{p}\cong 1+p\mathbb{Z}_{p},x\mapsto e^{px}$.
+And we notice that there is a continuous isomorphism
+$\mathbb{Z}_{p}$ $\cong 1+p\mathbb{Z}_{p}$
+$x$ $\mapsto e^{px}.$
 So the characters of $1+p\mathbb{Z}_{p}$ can be pulled back from the isomorphism as
-$\chi_{t}(x)=e^{2\pi i\{t\frac{\text{ln}x}{p}\}_{p}}$
-where $x\in 1+p\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$.
-Denoting $t=a_{-n}p^{-n}+a_{-n+1}p^{-n+1}+\dots+a_{-1}p^{-1}=\frac{a_{-n}+a_{-n+1}p+\dots+a_{-1}p^{n-1}}{p^{n}}$, and taking $c=a_{-n}+a_{-n+1}p+\dots+a_{-1}p^{n-1}$, we rewrite the character $\chi_{t}(x)$ as
-$\chi_{c,n+1}(x)=e^{2\pi i\{\frac{c\text{ln}x}{p^{n+1}}\}_{p}}$
-So $\forall x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$, the character of $\mathbb{Z}_{p}^{*}$ is trivial character $1$ and
-$\chi_{h,c,n}(x)=\chi_{h}(a_{0})\chi_{c,n}(\frac{x}{a_{0}})=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\{\frac{c\text{ln}\frac{x}{a_{0}}}{p^{n}}\}_{p}}$ (2.1)
-where $h\in\{0,1,\dots,p-2\},n\in\mathbb{N_{+}}$. And the conductor of $\chi_{h,c,n}$ is $n$.
+$\chi_{t}(x)=e^{2\pi i\left\{t\frac{\ln x}{p}\right\}_{p}},$
+where $x\in 1+p\mathbb{Z}_{p},t\in\mathbb{Q}_{p}/\mathbb{Z}_{p}$ and $\{\cdot\}_{p}$ denotes the $p$-adic fractional part. Letting
+$t=c_{-n}p^{-n}+c_{-n+1}p^{-n+1}+\dots+c_{-1}p^{-1}=\frac{c_{-n}+c_{-n+1}p+\dots+c_{-1}p^{n-1}}{p^{n}}$
+and taking
+$c=c_{-n}+c_{-n+1}p+\dots+c_{-1}p^{n-1},$
+we rewrite the character $\chi_{t}(x)$ as
+$\chi_{c,n+1}(x)=e^{2\pi i\left\{\frac{c\ln x}{p^{n+1}}\right\}_{p}}.$
+So $\forall x=a_{0}+a_{1}p+\dots\in\mathbb{Z}_{p}^{*}$ for $a_{0}\neq 0$, the characters of $\mathbb{Z}_{p}^{*}$ are trivial character $1$ and
+$\chi_{h,c,n}(x)=\chi_{h}(a_{0})\chi_{c,n}\left(\frac{x}{a_{0}}\right)=e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\frac{c\ln\left(\frac{x}{a_{0}}\right)}{p^{n}}\right\}_{p}},$ (2.1)
+where $h\in\{0,1,\dots,p-2\},n\in\mathbb{N_{+}}$. From (2.1) one can easily check that the conductor of the character $\chi_{h,c,n}$ is $n$.
 2.2 Eigenvalues and eigenfunctions for $D$ when $m=1$
@@
 $\langle\chi_{h,c,n},\chi_{h^{\prime},c^{\prime},n^{\prime}}\rangle$ $=\int_{\mathbb{Z}_{p}^{*}}\chi_{h,c,n}(x)\overline{\chi_{h^{\prime},c^{\prime},n^{\prime}}(x)}dx$
-$=\int_{\mathbb{Z}_{p}^{*}}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})\text{ln}\frac{x}{a_{0}}\}_{p}}dx$
-$=\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{1+p\mathbb{Z}_{p}}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})\text{ln}\frac{x}{a_{0}}\}_{p}}dx$
-$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{\mathbb{Z}_{p}}e^{2\pi i\{(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})y\}_{p}}dy$
-$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\boldsymbol{1}_{\mathbb{Z}_{p}}(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}})$
-$=(1-p^{-1})\delta_{h,c,n}^{h^{\prime},c^{\prime},n^{\prime}}$
+$=\int_{\mathbb{Z}_{p}^{*}}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)\ln\left(\frac{x}{a_{0}}\right)\right\}_{p}}dx$
+$=\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{1+p\mathbb{Z}_{p}}e^{2\pi i\left\{\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)\ln\left(\frac{x}{a_{0}}\right)\right\}_{p}}dx$
+$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\int_{\mathbb{Z}_{p}}e^{2\pi i\left\{\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)y\right\}_{p}}dy$
+$=p^{-1}\sum_{a_{0}=1}^{p-1}e^{\frac{2\pi i(h-h^{\prime})}{p-1}\text{ord}(a_{0})}\boldsymbol{1}_{\mathbb{Z}_{p}}\left(\frac{c}{p^{n}}-\frac{c^{\prime}}{p^{n^{\prime}}}\right)$
+$=(1-p^{-1})\delta_{h,c,n}^{h^{\prime},c^{\prime},n^{\prime}},$
 thus establishing orthogonality. ∎
@@
 Proof.
-Since the conductor of $\chi_{h,c,n}$ is $n$, it is trivial on $1+p^{n}\mathbb{Z}_{p}$ but nontrivial on $1+p^{n-1}\mathbb{Z}_{p}$.
-By the first orthogonality relation of the character, $\forall s\in\{1,2,\dots,n-1\}$, we have
-$\sum_{x\in(1+p^{s}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})}\chi_{h,c,n}(x)=0$
+[…]
+$\sum_{x\in(1+p^{s}\mathbb{Z}_{p})/(1+p^{n}\mathbb{Z}_{p})}\chi_{h,c,n}(x)=0,$
 so we only need to consider $z=a_{0}+a_{1}p+\dots+a_{n-1}p^{n-1}$ and $x=b_{0}+b_{1}p+\dots+b_{n-1}p^{n-1}$. Then we have
@@
 $=p^{-n}\sum_{j=0}^{n-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{n-1}\end{subarray}}p^{2j}(\chi_{h,c,n}(z)-\chi_{h,c,n}(x))$
-$=-p^{-n}((p-2)p^{n-1}+\sum_{d=1}^{n-2}(p-1)p^{n-1-d}p^{2d}+(p-1-(-1))p^{2(n-1)})\chi_{h,c,n}(x)$
-$=-p^{-n}((p-2)p^{n-1}+\sum_{d=1}^{n-2}(p-1)p^{n-1+d}+p^{2n-1})\chi_{h,c,n}(x)$
+$=-p^{-n}\left((p-2)p^{n-1}+\sum_{d=1}^{n-2}(p-1)p^{n-1-d}p^{2d}+(p-1-(-1))p^{2(n-1)}\right)\chi_{h,c,n}(x)$
+$=-p^{-n}\left((p-2)p^{n-1}+\sum_{d=1}^{n-2}(p-1)p^{n-1+d}+p^{2n-1}\right)\chi_{h,c,n}(x)$
 $=-p^{-n}(p^{2n-1}+p^{2n-2}-2p^{n-1})\chi_{h,c,n}(x)$
-$=-(p^{n-1}+p^{n-2}-\frac{2}{p})\chi_{h,c,n}(x)$
+$=-\left(p^{n-1}+p^{n-2}-\frac{2}{p}\right)\chi_{h,c,n}(x).$
 Hence $\chi_{h,c,n}$ is an eigenfunction of $D$ with eigenvalue $p^{n-1}-p^{n-2}+\frac{2}{p}$ for $n\geq 1$.
 Next, for any locally constant eigenfunction $f$ in $\mathbb{Z}_{p}^{*}$, we have the Fourier transform
-$\hat{f}(\xi)=\int_{\mathbb{Z}_{p}^{*}}f(x)\overline{\chi_{\xi}(x)}dx$
+$\hat{f}(\xi)=\int_{\mathbb{Z}_{p}^{*}}f(x)\overline{\chi_{\xi}(x)}dx,$
 then
-$f(x)=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x)$
+$f(x)=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x),$
 so we have
@@
 $=\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\lambda_{\xi}\chi_{\xi}(x)$
-$=\lambda_{f}\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x)$
+$=\lambda_{f}\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}\hat{f}(\xi)\chi_{\xi}(x),$
 thus
-$\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}(\lambda_{\xi}-\lambda_{f})\hat{f}(\xi)\chi_{\xi}(x)=0$
+$\sum_{\chi_{\xi}\in\text{Char}(\mathbb{Z}_{p}^{*})}(\lambda_{\xi}-\lambda_{f})\hat{f}(\xi)\chi_{\xi}(x)=0.$
 By orthogonality of eigenfunctions, we have $\lambda_{\xi}=\lambda_{f}$ if $\lambda_{\xi}$ belongs to the expansion of $f$.
@@
 Using the spectrum of $D$, we construct the Green’s function as
-$G(x,y)=\frac{1}{1-p^{-1}}\sum_{\phi_{h,c,n}\in\text{Char}(\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{h,c,n}(x)\overline{\phi_{h,c,n}(y)}}{\lambda_{n}}+C$ (2.2)
-Let $H_{n}$ denote the set of characters with conductor at most $n$ and define $C_{n}:=H_{n}\backslash H_{n-1}$. By the second orthogonality relation of the character, we have
-$\sum_{\phi\in H_{n}}\phi(x)\overline{\phi(y)}=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})|H_{n}|$ (2.3)
+$G(x,y)=\frac{1}{1-p^{-1}}\sum_{\phi_{h,c,n}\in\text{Char}(\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{h,c,n}(x)\overline{\phi_{h,c,n}(y)}}{\lambda_{n}}+C.$ (2.2)
+Letting $H_{n}$ denote the set of characters with conductor at most $n$ and defining $C_{n}:=H_{n}\backslash H_{n-1}$. By the second orthogonality relation of the characters, we have
+$\sum_{\phi\in H_{n}}\phi(x)\overline{\phi(y)}=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})|H_{n}|,$ (2.3)
 where $|H_{n}|=(p-1)p^{n-1}=\varphi(p^{n})$.
-And we notice that if $v_{p}(x-y)=M$ and $n\geq M+1$, the sum (2.3) vanishes.
+If $v_{p}(x-y)=M$ and $n\geq M+1$, the sum (2.3) vanishes.
 [^1] The idea of the finite sum formula for the case $m=1$ originated from Ekam’s independent study.
@@
 $G(x,y)$ $=\frac{1}{1-p^{-1}}\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n}}+C$ (2.4)
-$=\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n}}-\frac{\varphi(p^{M})}{\lambda_{M+1}})+C$
-$=-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-2}+\frac{p^{M+1}}{p^{M+1}+p^{M}-2}+C$
+$=\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n}}-\frac{\varphi(p^{M})}{\lambda_{M+1}}\right)+C$
+$=-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-2}+\frac{p^{M+1}}{p^{M+1}+p^{M}-2}+C,$
 which is expressed as a finite sum.
@@
 $D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)$ (2.5)
-$=$ $|x|_{p}\int_{\cup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$
-$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz-\left(|x|_{p}\int_{\cup_{k\neq l}p^{k}\mathbb{Z}_{p}^{*}}\frac{1}{|z-x|_{p}^{2}}dz\right)\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
+$=$ $|x|_{p}\int_{\bigcup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$
+$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz-\left(|x|_{p}\int_{\bigcup_{k\neq l}p^{k}\mathbb{Z}_{p}^{*}}\frac{1}{|z-x|_{p}^{2}}dz\right)\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
 where the first term is
 $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-x|_{p}^{2}}dz$ (2.6)
-$=$ $|x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{p^{-2l}|p^{-l}z-p^{-l}x|_{p}^{2}}dz$
 $=$ $|p^{-l}x|_{p}\int_{p^{l}\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(p^{-l}z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|p^{-l}z-p^{-l}x|_{p}^{2}}dp^{-l}z$
-$=$ $|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz$
-when $v_{p}(x)=l$, (LABEL:S:eq_3) reduces to
-$\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)}{|z-p^{-l}x|_{p}^{2}}dz=(D_{0}\phi_{h,c,n})(p^{-l}x)=\lambda_{n}\phi_{h,c,n}(p^{-l}x)$
+$=$ $|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz,$
+when $v_{p}(x)=l$, (LABEL:eq:compute_eigfunction_general_m_first_term) reduces to
+$\int_{\mathbb{Z}_{p}^{*}}\frac{\phi_{h,c,n}(z)-\phi_{h,c,n}(p^{-l}x)}{|z-p^{-l}x|_{p}^{2}}dz=(D_{0}\phi_{h,c,n})(p^{-l}x)=\lambda_{n}\phi_{h,c,n}(p^{-l}x),$
 while the second term is
-$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}}{\text{max}\{p^{-k},p^{-l}\}^{2}}p^{-k}(1-p^{-1})\phi_{h,c,n}(p^{-l}x)$
+$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}}{\max\{p^{-k},p^{-l}\}^{2}}p^{-k}(1-p^{-1})\phi_{h,c,n}(p^{-l}x).$
 Then the coefficient of $\phi_{h,c,n}(p^{-l}x)$ simplifies to
-$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}p^{-k}}{\text{max}\{p^{-k},p^{-l}\}^{2}}(1-p^{-1})$
-$=$ $-(1-p^{-1})(\sum_{k=0}^{l-1}p^{-l}p^{-k}p^{2k}+\sum_{k=l+1}^{m-1}p^{-l}p^{-k}p^{2l})$
-$=$ $-(1-p^{-1})(p^{-l}\sum_{k=0}^{l-1}p^{k}+\sum_{k=l+1}^{m-1}p^{-(k-l)})$
-$=$ $\frac{p^{-l}+p^{l+1-m}-2}{p}$
-thus (LABEL:S:eq_2) becomes
+$-\sum_{\begin{subarray}{c}k=0\\ k\neq l\end{subarray}}^{m-1}\frac{p^{-l}p^{-k}}{\max\{p^{-k},p^{-l}\}^{2}}(1-p^{-1})$
+$=$ $-(1-p^{-1})\left(\sum_{k=0}^{l-1}p^{-l}p^{-k}p^{2k}+\sum_{k=l+1}^{m-1}p^{-l}p^{-k}p^{2l}\right)$
+$=$ $-(1-p^{-1})\left(p^{-l}\sum_{k=0}^{l-1}p^{k}+\sum_{k=l+1}^{m-1}p^{-(k-l)}\right)$
+$=$ $\frac{p^{-l}+p^{l+1-m}-2}{p},$
+thus (LABEL:eq:compute_eigfunction_general_m) becomes
 $-(p^{n-1}+p^{n-2}-\frac{2}{p}-\frac{p^{-l}+p^{l+1-m}-2}{p})\phi_{h,c,n}(p^{-l}x)$
-$=$ $-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})\phi_{h,c,n}(p^{-l}x)$
-when $v_{p}(x)\neq l$, (LABEL:S:eq_3) becomes
-$\frac{|p^{-l}x|_{p}}{\text{max}\{1,|p^{-l}x|_{p}\}^{2}}\int_{\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(z)dz=0$
+$=$ $-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})\phi_{h,c,n}(p^{-l}x).$
+When $v_{p}(x)\neq l$, (LABEL:eq:compute_eigfunction_general_m_first_term) becomes
+$\frac{|p^{-l}x|_{p}}{\max\{1,|p^{-l}x|_{p}\}^{2}}\int_{\mathbb{Z}_{p}^{*}}\phi_{h,c,n}(z)dz=0$
 since $\phi_{h,c,n}$ is orthogonal to $\phi_{0}=1$ by Lemma 2.1, and the second term vanishes. Denoting $\lambda_{n,l}=-p^{n-1}-p^{n-2}+\frac{p^{-l}+p^{l+1-m}}{p}$, we obtain
-$D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)=\lambda_{n,l}\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
+$D(\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x))(x)=\lambda_{n,l}\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
 which means $\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$ is the eigenfunction of $D$ with the eigenvalue $\lambda_{n,l}$.
 When $n=0$, we consider $\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$, similarly, we have
-$D\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{1-\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz+\frac{p^{-l}+p^{l+1-m}-2}{p}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)$
+$D\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=|p^{-l}x|_{p}\int_{\mathbb{Z}_{p}^{*}}\frac{1-\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)}{|z-p^{-l}x|_{p}^{2}}dz+\frac{p^{-l}+p^{l+1-m}-2}{p}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
 so when $v_{p}(x)=l$, the first part vanishes, and when $v_{p}(x)\neq l$, the first part becomes
-$\frac{|p^{-l}x|_{p}(1-p^{-1})}{\text{max}\{1,|p^{-l}x|_{p}\}^{2}}=(1-p^{-1})p^{-|v_{p}(x)-l|}$
+$\frac{|p^{-l}x|_{p}(1-p^{-1})}{\max\{1,|p^{-l}x|_{p}\}^{2}}=(1-p^{-1})p^{-|v_{p}(x)-l|},$
 denoting this vector by $\boldsymbol{r}_{l}=\begin{pmatrix}r_{0l}\\ r_{1l}\\ \vdots\\ r_{m-1,l}\end{pmatrix}$.
 We want to use the linear combination of them to construct the eigenvalue and eigenfunction:
-$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\sum_{l=0}^{m-1}c_{l}\boldsymbol{r}_{l}=(\boldsymbol{r}_{0},\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{m-1})\begin{pmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{m-1}\end{pmatrix}=\boldsymbol{R}\boldsymbol{c}$
+$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\sum_{l=0}^{m-1}c_{l}\boldsymbol{r}_{l}=(\boldsymbol{r}_{0},\boldsymbol{r}_{1},\dots,\boldsymbol{r}_{m-1})\begin{pmatrix}c_{0}\\ c_{1}\\ \vdots\\ c_{m-1}\end{pmatrix}=\boldsymbol{R}\boldsymbol{c},$
 where we require that
-$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\boldsymbol{c}$
+$D\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\sum_{l=0}^{m-1}c_{l}\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x)=\lambda\boldsymbol{c},$
 so we only need to solve
-$\boldsymbol{R}\boldsymbol{c}=\lambda\boldsymbol{c}$
+$\boldsymbol{R}\boldsymbol{c}=\lambda\boldsymbol{c},$
 […]
+$\phi_{0}^{l}(x)=\boldsymbol{q}_{l}(x)=\sum_{k=0}^{m-1}q_{k,l}\boldsymbol{1}_{p^{k}\mathbb{Z}_{p}^{*}}(x)$
+and
+$\phi_{h,c,n}^{l}(x)=\phi_{h,c,n}(p^{-l}x)\boldsymbol{1}_{p^{l}\mathbb{Z}_{p}^{*}}(x),$
+we have the following lemma:
 Lemma 2.3.
@@
 $=\delta_{l}^{l^{\prime}}\langle\phi_{h,c,n},\phi_{h^{\prime},c^{\prime},n^{\prime}}\rangle$
-$=(1-p^{-1})\delta_{h,c,n,l}^{h^{\prime},c^{\prime},n^{\prime},l^{\prime}}$
-which is also orthogonal. Lastly, $\phi_{h,c,n}^{l},\phi_{0}^{l^{\prime}}>=0$ is trivial as $\phi_{h,c,n}^{l}$ is not a constant function.
-For any eigenfunction $f$, we have the decomposition $f=f_{c}+f_{0}+f_{1}+\dots+f_{m-1}$, where $f_{c}$ is constant on each $p^{k}\mathbb{Z}_{p}^{*}$ and $\text{supp}f_{l}\subseteq p^{l}\mathbb{Z}_{p}^{*}$ with zero DC component (i.e.
+$=(1-p^{-1})\delta_{h,c,n,l}^{h^{\prime},c^{\prime},n^{\prime},l^{\prime}},$
+which is also orthogonal. Lastly, $\langle\phi_{h,c,n}^{l},\phi_{0}^{l^{\prime}}\rangle=0$ is trivial as $\phi_{h,c,n}^{l}$ is not a constant function on each stratum $p^{k}\mathbb{Z}_{p}^{*}$.
+For any eigenfunction $f$, we have the decomposition $f=f_{c}+f_{0}+f_{1}+\dots+f_{m-1}$, where $f_{c}$ is constant on each stratum and $\text{supp}f_{l}\subseteq p^{l}\mathbb{Z}_{p}^{*}$ with zero DC component (i.e.
 the Fourier expansion of $f_{l}$ don’t have trivial character).
@@
 $\lambda_{n,l}$ may not be disjoint as $\lambda_{n,l}=\lambda_{n,m-l}$ for $0\leq l\leq m-1,n\geq 1$.
-Then we have the Green’s function as
-$G(x,y)=\frac{1}{1-p^{-1}}(\sum_{k=0}^{m-1}\sum_{\phi_{n}^{k}\in\text{Char}(p^{k}\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{n}^{k}(x)\overline{\phi_{n}^{k}(y)}}{\lambda_{n,k}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}})+C(|y|_{p})$ (2.7)
-Similarly, let $H_{n}^{l}$ denote the set of characters generated from $p^{l}\mathbb{Z}_{p}^{*}$ with conductor at most $n$ and define $C_{n}^{l}:=H_{n}^{l}\backslash H_{n-1}^{l}$.
-Consider $\sum_{k=0}^{m-1}\displaystyle\sum_{\phi^{k}\in H_{n}^{k}}\phi^{k}(x)\overline{\phi^{k}(y)}$.
-When $v_{p}(x)\neq v_{p}(y)$ (i.e.
-$v_{p}(d(x,y))=0$), the sum vanishes.
-When $v_{p}(x)=v_{p}(y)=l$ (i.e.
-$v_{p}(d(x,y))>0$), then
+Then the Green’s function is
+$G(x,y)=\frac{1}{1-p^{-1}}\left(\sum_{k=0}^{m-1}\sum_{\phi_{n}^{k}\in\text{Char}(p^{k}\mathbb{Z}_{p}^{*})\backslash\{1\}}\frac{\phi_{n}^{k}(x)\overline{\phi_{n}^{k}(y)}}{\lambda_{n,k}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}}\right)+C(|y|_{p}).$ (2.7)
+Similarly, letting $H_{n}^{l}$ denote the set of characters generated from $p^{l}\mathbb{Z}_{p}^{*}$ with conductor at most $n$ and defining $C_{n}^{l}:=H_{n}^{l}\backslash H_{n-1}^{l}$. Considering
+$\sum_{k=0}^{m-1}\sum_{\phi^{k}\in H_{n}^{k}}\phi^{k}(x)\overline{\phi^{k}(y)}.$
+When $v_{p}(x)\neq v_{p}(y)$ (i.e. $v_{p}(d(x,y))=0$), the sum vanishes. When $v_{p}(x)=v_{p}(y)=l$ (i.e. $v_{p}(d(x,y))\geq 0$), then
 $\sum_{\phi^{l}\in H_{n}^{l}}\phi^{l}(x)\overline{\phi^{l}(y)}$ $=\sum_{\phi^{0}\in H_{n}^{0}}\phi^{0}(p^{-l}x)\overline{\phi^{0}(p^{-l}y)}$
-$=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})\varphi(p^{n})$
+$=\boldsymbol{1}_{1+p^{n}\mathbb{Z}_{p}}(xy^{-1})\varphi(p^{n}).$
 So when $v_{p}(d(x,y))=M$ and $v_{p}(x)=v_{p}(y)=l$, we have the Green’s function
-$G(x,y)=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n,l}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}})+C(|y|_{p})$ (2.8)
-$=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n,l}}-\frac{\varphi(p^{M})}{\lambda_{M+1,l}}+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{\lambda_{0,k}})+C(|y|_{p})$
+$G(x,y)=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}\frac{\phi(x)\overline{\phi(y)}}{\lambda_{n,l}}+\sum_{k=1}^{m-1}\frac{\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}}{\lambda_{0,k}}\right)+C(|y|_{p})$ (2.8)
+$=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M}\frac{\varphi(p^{n})-\varphi(p^{n-1})}{\lambda_{n,l}}-\frac{\varphi(p^{M})}{\lambda_{M+1,l}}+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{\lambda_{0,k}}\right)+C(|y|_{p})$
 $=$ $-\sum_{n=1}^{M}\frac{(p-1)p^{n}}{p^{n}+p^{n-1}-(p^{-l}+p^{l+1-m})}+\frac{p^{M+1}}{p^{M+1}+p^{M}-(p^{-l}+p^{l+1-m})}$
-$+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{(1-p^{-1})\lambda_{0,k}}+C(|y|_{p})$
+$+\sum_{k=1}^{m-1}\frac{q_{v_{p}(x)k}q_{v_{p}(y)k}}{(1-p^{-1})\lambda_{0,k}}+C(|y|_{p}).$
 Finally, we verify that $G(x,y)$ is actually the Green’s function:
@@
 […]
-$=\frac{1}{1-p^{-1}}((\sum_{n=1}^{M}(\varphi(p^{n})-\varphi(p^{n-1}))-\varphi(p^{M}))\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=1}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k})$
-$=\frac{1}{1-p^{-1}}(-\varphi(1)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=0}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}-q_{v_{p}(x)0}q_{v_{p}(y)0})$
-$=\frac{1}{1-p^{-1}}(-\delta_{v_{p}(x)}^{v_{p}(y)}+\delta_{v_{p}(x)}^{v_{p}(y)}-\frac{1}{m})$
+$=\frac{1}{1-p^{-1}}\left(\left(\sum_{n=1}^{M}(\varphi(p^{n})-\varphi(p^{n-1}))-\varphi(p^{M})\right)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=1}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}\right)$
+$=\frac{1}{1-p^{-1}}\left(-\varphi(1)\delta_{v_{p}(x)}^{v_{p}(y)}+\sum_{k=0}^{m-1}q_{v_{p}(x)k}q_{v_{p}(y)k}-q_{v_{p}(x)0}q_{v_{p}(y)0}\right)$
+$=\frac{1}{1-p^{-1}}\left(-\delta_{v_{p}(x)}^{v_{p}(y)}+\delta_{v_{p}(x)}^{v_{p}(y)}-\frac{1}{m}\right)$
 $=-\frac{1}{m(1-p^{-1})}$
@@
 Then $DG(x,y)+\frac{1}{m(1-p^{-1})}$ is a distribution supported on the diagonal. And we notice that
-$\int_{\cup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}DG(x,y)+\frac{1}{m(1-p^{-1})}dx$ (2.10)
+$\int_{\bigcup_{k=0}^{m-1}p^{k}\mathbb{Z}_{p}^{*}}DG(x,y)+\frac{1}{m(1-p^{-1})}dx$
 […]
-$=$ $1$
+$=$ $1.$
 So $G(x,y)$ satisfies $DG(x,y)=\delta_{y}(x)-\frac{1}{m(1-p^{-1})}$.
-We can compute the heat kernel by using spectrum of $D$ as
-$H(t,x,y)=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}e^{\lambda_{n,l}t}\phi(x)\overline{\phi(y)}+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)})$ (2.11)
-$=$ $\frac{1}{1-p^{-1}}(\sum_{n=1}^{M}e^{\lambda_{n,l}t}(\varphi(p^{n})-\varphi(p^{n-1}))-e^{\lambda_{M+1,l}t}\varphi(p^{M})+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k})$
+Furthermore, we can compute the heat kernel by using spectrum of $D$ as
+$H(t,x,y)=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M+1}\sum_{\phi\in C_{n}^{l}}e^{\lambda_{n,l}t}\phi(x)\overline{\phi(y)}+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}\phi_{0}^{k}(x)\overline{\phi_{0}^{k}(y)}\right)$ (2.9)
+$=$ $\frac{1}{1-p^{-1}}\left(\sum_{n=1}^{M}e^{\lambda_{n,l}t}(\varphi(p^{n})-\varphi(p^{n-1}))-e^{\lambda_{M+1,l}t}\varphi(p^{M})+\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k}\right)$
 $=$ $(p-1)\sum_{n=1}^{M}e^{-(p^{n-1}+p^{n-2}-\frac{p^{-l}+p^{l+1-m}}{p})t}p^{n-1}-e^{-(p^{M}+p^{M-1}-\frac{p^{-l}+p^{l+1-m}}{p})t}p^{M}$
-$+\frac{p}{p-1}\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k}$
+$+\frac{p}{p-1}\sum_{k=1}^{m-1}e^{\lambda_{0,k}t}q_{v_{p}(x)k}q_{v_{p}(y)k},$
 which is also an important fundamental quantity.
+Remark 2.6.
+As $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}\cong p^{\mathbb{Z}/m\mathbb{Z}}\times\mathbb{Z}_{p}^{*}$ is a multiplicative group, and the characters of $p^{\mathbb{Z}/m\mathbb{Z}}$ is given by
+$\chi_{k}(p^{l})=e^{\frac{2\pi ik}{m}l}$
+for $k,l\in\{0,1,\dots,m-1\}$. So $\forall\ x=p^{l}(a_{0}+a_{1}p+\dots)\in\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with $a_{0}\neq 0$, we have the character
+$\chi_{k,h,c,n}(x)=\chi_{k}(p^{l})\chi_{h}(a_{0})\chi_{c,n}\left(\frac{x}{p^{l}a_{0}}\right)=e^{\frac{2\pi ik}{m}l}e^{\frac{2\pi ih}{p-1}\text{ord}(a_{0})}e^{2\pi i\left\{\frac{c\ln\left(\frac{x}{p^{l}a_{0}}\right)}{p^{n}}\right\}_{p}}.$
+One can check that the eigenfunctions are the linear transformation of the characters on each stratum.
+So the orthogonality and completeness can be easily obtained by the orthogonality of characters and Fourier transform.
+Remark 2.7.
+For general $s$ case, we define the operator $D^{s}$ on the Tate curve as
+$D^{s}u(x)=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/q^{\mathbb{Z}}}\frac{u(z)-u(x)}{|z-x|_{p}^{1+s}}dz,$
+which is also self-adjoint, negative semi-definite and preserves locally constant functions.
+One can check that all our computations carry over to $D^{s}$, simply by replacing $2$ with $1+s$, so that the eigenfunctions of $D^{s}$ are the same with $D$.
+Hence, we may use the same method to determine the spectrum of $D^{s}$, and then obtain its Green’s function and heat kernel as a finite sum formula.
 3 Mean field equation on the finite quotient
 In this section, we consider the mean field equation on the Tate curve
-$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ (3.1)
+$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}.$ (3.1)
 To study this equation, we first consider this equation on a finite quotient
-$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ (3.2)
+$Du+\rho e^{u}=\rho\delta_{y}\text{ in }\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p}).$ (3.2)
 i.e.
@@
 $Du(x)$ $=|x|_{p}\int_{\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})}|z|_{p}\frac{u(z)-u(x)}{|z-x|_{p}^{2}}p^{-d}d\delta$ (3.3)
-$=p^{-d}\sum_{\begin{subarray}{c}0\leq k\leq m-1\\ 1\leq a_{0}\leq p-1\\ 0\leq a_{i}\leq p-1\\ 1\leq i\leq d-1\end{subarray}}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x))$
+$=p^{-d}\sum_{\begin{subarray}{c}0\leq k\leq m-1\\ 1\leq a_{0}\leq p-1\\ 0\leq a_{i}\leq p-1\\ 1\leq i\leq d-1\end{subarray}}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x)).$
 So the flat Laplacian on the quotient space of the Tate curve reduces to a graph Laplacian on a finite weighted graph.
@@
 Then (3.2) can be rewritten as a matrix equation
-$\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\rho p^{d}\boldsymbol{e}_{y}$ (3.4)
+$\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\rho p^{d}\boldsymbol{e}_{y}.$ (3.4)
 From (3.3), $\boldsymbol{A}\in\mathbb{R}^{m(p-1)p^{d-1}\times m(p-1)p^{d-1}}$ is a symmetric matrix with zero row and colum sums.
@@
 3.1 Structure of the solution
-In this part, we present structural properties of the solution to (3.4). These properties are observed from numerical results.
 Lemma 3.2.
-If $\boldsymbol{u}$ solves (3.4), then $u_{i}\leq d\text{ln}p,\forall i$. The equality holds if and only if $p=2,m=1$.
+If $\boldsymbol{u}$ solves (3.4), then $u_{i}\leq d\ln p,\forall i$. The equality holds if and only if $p=2,m=1$.
 Proof.
 From (3.4), we have
-$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=\rho p^{d}\delta_{iy},\ \forall i$ (3.5)
+$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=\rho p^{d}\delta_{iy},\ \forall i.$ (3.5)
 Summing over $i$ yields
-$\sum_{i,j}a_{ij}u_{j}+\rho\sum_{i}e^{u_{i}}=\rho p^{d}$
+$\sum_{i,j}a_{ij}u_{j}+\rho\sum_{i}e^{u_{i}}=\rho p^{d},$
 since $\sum_{i,j}a_{ij}u_{j}=\sum_{j}u_{j}\sum_{i}a_{ij}=0$, we obtain
-$\sum_{i}e^{u_{i}}=p^{d}$
-then $e^{u_{i}}\leq p^{d}$, thus $u_{i}\leq d\text{ln}p$ for all $i$. And the equality holds if and only if $p=2,m=1$ as there is one term in the sum. ∎
+$\sum_{i}e^{u_{i}}=p^{d},$
+then $e^{u_{i}}\leq p^{d}$, thus $u_{i}\leq d\ln p$ for all $i$. And the equality holds if and only if $p=2,m=1$ as there only has one term in the sum. ∎
 Lemma 3.3.
-$u_{y}$ is the only smallest in $\boldsymbol{u}$.
+$u_{y}$ is the only smallest one in $\boldsymbol{u}$.
 Proof.
-Let $m=\text{min}_{i}\ u_{i}$. Suppose $\exists\ i\neq y$, s.t. $u_{i}=m$. From (3.5) with $i\neq y$,
-$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=0$
+Let $m=\min_{i}u_{i}$. Suppose $\exists\ i\neq y$, s.t. $u_{i}=m$. From (3.5) with $i\neq y$,
+$\sum_{j}a_{ij}u_{j}+\rho e^{u_{i}}=0,$
 then $\rho e^{u_{i}}+a_{ii}u_{i}=-\sum_{j\neq i}a_{ij}u_{j}\leq-m\sum_{j\neq i}a_{ij}=ma_{ii}$, so we have
-$\rho e^{u_{i}}\leq a_{ii}(u_{i}-m)=0$
+$\rho e^{u_{i}}\leq a_{ii}(u_{i}-m)=0,$
 a contradiction. ∎
@@
 ∎
-Let $\boldsymbol{u}=\begin{pmatrix}\boldsymbol{u}_{1}\\ \boldsymbol{u}_{2}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, $\boldsymbol{\sigma}=\begin{pmatrix}&&&1\\ 1&&&\\ &\ddots&&\\ &&1&\end{pmatrix}\in\mathbb{R}^{(p-1)p^{d-1}\times(p-1)p^{d-1}}$ be the shift matrix on blocks of size $(p-1)p^{d-1}$, and define $\sigma_{k}(\boldsymbol{u})=\begin{pmatrix}\boldsymbol{u}_{1}\\ \vdots\\ \boldsymbol{\sigma}\boldsymbol{u}_{k}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}$.
-When $i\neq j$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ as $\boldsymbol{B}_{ij}=p^{-|i-j|}\boldsymbol{1}\boldsymbol{1}^{\top}$ when $i\neq j$.
+Let $\boldsymbol{u}=\begin{pmatrix}\boldsymbol{u}_{1}\\ \boldsymbol{u}_{2}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$ where $\boldsymbol{u}_{i}\in\mathbb{R}^{(p-1)p^{d-1}}$, $\boldsymbol{\sigma}=\begin{pmatrix}&&&1\\ 1&&&\\ &\ddots&&\\ &&1&\end{pmatrix}\in\mathbb{R}^{(p-1)p^{d-1}\times(p-1)p^{d-1}}$ be the shift matrix on blocks of size $(p-1)p^{d-1}$, and define $\sigma_{k}(\boldsymbol{u})=\begin{pmatrix}\boldsymbol{u}_{1}\\ \vdots\\ \boldsymbol{\sigma}\boldsymbol{u}_{k}\\ \vdots\\ \boldsymbol{u}_{m}\end{pmatrix}$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}$ as $\boldsymbol{A}_{ij}$ is circulant.
+Furthermore, when $i\neq j$, then $\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ as $\boldsymbol{B}_{ij}=p^{-|i-j|}\boldsymbol{1}\boldsymbol{1}^{\top}$ when $i\neq j$.
+Therefore, we have
+$\boldsymbol{\sigma}\boldsymbol{A}_{ii}=\boldsymbol{A}_{ii}\boldsymbol{\sigma}$ (3.6)
+and
+$\boldsymbol{\sigma}\boldsymbol{A}_{ij}=\boldsymbol{A}_{ij}\boldsymbol{\sigma}=\boldsymbol{A}_{ij}$ (3.7)
+when $i\neq j$.
+Remark 3.5.
+[…]
 Lemma 3.6.
@@
 As $\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}=\boldsymbol{b}$, so we have
-$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{k}}=\boldsymbol{b}_{k}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
+$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{k}}=\boldsymbol{b}_{k},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}.\end{cases}$
 Multiplying the $k$-th block by $\boldsymbol{\sigma}$ gives
-$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{\sigma}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho\boldsymbol{\sigma}e^{\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k}\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
+$\begin{cases}\displaystyle\sum_{j}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{\sigma}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\rho\boldsymbol{\sigma}e^{\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k},\\ \qquad\qquad\quad\vdots\\ \displaystyle\sum_{j}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m},\end{cases}$
 which is equivalent to
-$\begin{cases}\displaystyle\sum_{j\neq k}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\boldsymbol{A}_{1k}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1}\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\boldsymbol{A}_{kk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{\sigma}\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k}\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\boldsymbol{A}_{mk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}\end{cases}$
-Thus $\sigma_{k}(\boldsymbol{u})$ is a solution of $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\sigma_{k}(\boldsymbol{b})$. ∎
-Lemma 3.4 and Lemma 3.5 imply certain symmetric properties: given one solution, we can construct many solutions by folding or rotating the solution.
-However, if the solution is radial symmetric, then Lemma 3.4 and Lemma 3.5 are trivial.
-From numerical results we can see that there indeed have radial symmetric solutions, and by the following theorem we proved that there only exist radial symmetric solutions when $\rho$ is small.
+$\begin{cases}\displaystyle\sum_{j\neq k}\boldsymbol{A}_{1j}\boldsymbol{u}_{j}+\boldsymbol{A}_{1k}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{1}}=\boldsymbol{b}_{1},\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{kj}\boldsymbol{u}_{j}+\boldsymbol{A}_{kk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{\sigma}\boldsymbol{u}_{k}}=\boldsymbol{\sigma}\boldsymbol{b}_{k},\\ \qquad\qquad\qquad\qquad\vdots\\ \displaystyle\sum_{j\neq k}\boldsymbol{A}_{mj}\boldsymbol{u}_{j}+\boldsymbol{A}_{mk}\boldsymbol{\sigma}\boldsymbol{u}_{k}+\rho e^{\boldsymbol{u}_{m}}=\boldsymbol{b}_{m}.\end{cases}$
+by (3.6) and (3.7). Thus $\sigma_{k}(\boldsymbol{u})$ is a solution of $\boldsymbol{A}\boldsymbol{v}+\rho e^{\boldsymbol{v}}=\sigma_{k}(\boldsymbol{b})$. ∎
+Lemma 3.4 and Lemma 3.6 imply certain symmetric properties: since $\boldsymbol{b}=p^{d}\boldsymbol{e}_{y}$, so given one solution, we can construct many solutions by folding or rotating the solution unless the solution is radially symmetric.
+And by the following theorem we proved that there only exist radially symmetric solutions when $\rho$ is small.
 Theorem 3.7.
-When $\rho\leq p^{-d}-p^{-d-m}$, denote $u_{y}(x)$ be the solution of (3.4), then $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$.
-Moreover, if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ and $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$, which means that $u(x)$ is radial symmetric.
+When $\rho\leq p^{-d}-p^{-d-m}$, denote $u_{y}(x)$ be the solution of (3.4), then $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, which means that the solution is radially symmetric.
+Moreover, if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ and $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$.
 Proof.
-First, from (3.3), we compute
+From (3.3), we compute
 $p^{d}Du(x)=$ $\sum_{z\neq x}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}(u(z)-u(x))$ (3.8)
 $=$ $\sum_{z\neq x}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}u(z)-\sum_{z\neq x}\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}u(x)$
-$=$ $\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
-$-(\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j})u(x)$
+$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
+$-\left(\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}\right)u(x).$
 Since $a_{0}\neq b_{0}$ correspond to $(p-2)p^{d-1}$ cases and $a_{i}=b_{i},\ i=0,1,\dots,j-1,\ a_{j}\neq b_{j}$ has $(p-1)p^{d-1-j}$ cases for $1\leq j\leq d-1$, so the coefficient of $u(x)$ is
-$\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}$ (3.7)
-$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d}\sum_{i=0}^{d-2}p^{i}$
+$\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}$ (3.9)
+$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d-1-j}\sum_{j=1}^{d-1}p^{2j}$
+$=$ $(p-1)p^{d-1}\sum_{v_{p}(z)\neq v_{p}(x)}p^{-|v_{p}(z)-v_{p}(x)|}+(p-2)p^{d-1}+(p-1)p^{d}\sum_{j=0}^{d-2}p^{j}$
 $=$ $(p-1)p^{d-1}\frac{2-|x|_{p}-p^{1-m}|x|_{p}^{-1}}{p-1}+(p-2)p^{d-1}+p^{d}(p^{d-1}-1)$
 $=$ $(2-|x|_{p}-p^{1-m}|x|_{p}^{-1})p^{d-1}+p^{2d-1}-2p^{d-1}$
-$=$ $p^{2d-1}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1}$
-which coincides with the negative of the diagonal entries of $\boldsymbol{B}$. Note that
+$=$ $p^{2d-1}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1},$
+which coincides with the negative of the diagonal entries of $\boldsymbol{B}$.
+Note that
 $\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{2(d-1)}u(x)$
-$=$ $\sum_{\begin{subarray}{c}a_{0}\neq b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}p^{2}u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\neq b_{d-1}\end{subarray}}p^{2(d-1)}u(z)$
-$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-1)u(z)$
-$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}=b_{1}\\ a_{2}\neq b_{2}\\ a_{3}\dots a_{d-1}\end{subarray}}(p^{4}-p^{2})u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2})u(z)$
+$=$ $\sum_{\begin{subarray}{c}a_{0}\neq b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}p^{2}u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}p^{2(d-1)}u(z)$
+$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\neq b_{1}\\ a_{2}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-1)u(z)$
+$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}=b_{1}\\ a_{2}\neq b_{2}\\ a_{3}\dots a_{d-1}\end{subarray}}(p^{4}-p^{2})u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2})u(z)$
 $=$ $\cdots$
-$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2(d-2)})u(z)$
-$=$ $\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$
+$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ a_{1}\dots a_{d-1}\end{subarray}}(p^{2}-1)u(z)+\dots+\sum_{\begin{subarray}{c}a_{0}=b_{0}\\ \vdots\\ a_{d-2}=b_{d-2}\\ a_{d-1}\end{subarray}}(p^{2(d-1)}-p^{2(d-2)})u(z)$
+$=$ $\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z),$
 so we have
-$p^{d}Du(x)=$ $\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)$
-$+\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$
-$-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)$
+$p^{d}Du(x)=$ $\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)$
+$+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x),$
 then when $x\neq y$ the equation (3.2) becomes
-$\sum_{v_{p}(z)=0}^{v_{p}(x)-1}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)+1}^{m-1}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{v_{p}(z)=v_{p}(x)}u(z)+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)$ (3.8)
-$+\rho p^{d}e^{u(x)}-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)=0$
-Denote $u(x)=u_{l}(b_{0},b_{1},\dots,b_{d-1})$, by fixing $b_{0},b_{1},\dots,b_{d-2}$ and varying $b_{d-1}$, we see that $u(x)$ satisfies (LABEL:1_4).
-In addition, the function $f(x)=\rho p^{d}e^{x}-kx,\ x\leq d\text{ln}p$ is monotonic when $p^{2d}\rho\leq k$.
+$\rho p^{d}e^{u(x)}+\sum_{\begin{subarray}{c}z\\ 0\leq v_{p}(z)\leq v_{p}(x)-1\end{subarray}}\frac{|x|_{p}}{|z|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(x)+1\leq v_{p}(z)\leq m-1\end{subarray}}\frac{|z|_{p}}{|x|_{p}}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=v_{p}(x)\end{subarray}}u(z)$ (3.10)
+$+\sum_{j=0}^{d-2}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u(z)-(p^{2d-1}+p^{2d-2}-(|x|_{p}+p^{1-m}|x|_{p}^{-1})p^{d-1})u(x)=0.$
+Denoting $u(x)=u_{l}(b_{0},b_{1},\dots,b_{d-1})$, by fixing $b_{0},b_{1},\dots,b_{d-2}$ and varying $b_{d-1}$, we see that $u(x)$ satisfies (LABEL:eq:equation_for_x_neq_y).
+In addition, the function $f(x)=\rho p^{d}e^{x}-kx,\ x\leq d\ln p$ is monotone when $p^{2d}\rho\leq k$.
 Then by Lemma 3.2, if $p^{2d}\rho\leq p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-1})$ is constant for all $b_{d-1}$ except at $x=y$.
 Hence for $x\neq y$, we denote $u(x)$ by $u_{l}(b_{0},b_{1},\dots,b_{d-2})$.
-When $v_{p}(x)=l\neq v_{p}(y)$, (LABEL:1_4) becomes
-$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{v_{p}(z)=l}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$
-$+p(p^{2(d-1)}-p^{2(d-2)})u_{l}(b_{0},b_{1},\dots,b_{d-2})+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}$
-$-(p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0$
+When $v_{p}(x)=l\neq v_{p}(y)$, (LABEL:eq:equation_for_x_neq_y) becomes
+$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=l\end{subarray}}u(z)$
+$+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})+p(p^{2(d-1)}-p^{2(d-2)})u_{l}(b_{0},b_{1},\dots,b_{d-2})$
+$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-1}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0.$
 This simplifies to
-$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{v_{p}(z)=l}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$ (3.9)
-$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0$
-Fixing $b_{0},b_{1},\dots,b_{d-3}$ and varying $b_{d-2}$, we see that $u(x)$ satisfies (LABEL:1_5).
+$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=l\end{subarray}}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$ (3.11)
+$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0.$
+Fixing $b_{0},b_{1},\dots,b_{d-3}$ and varying $b_{d-2}$, we see that $u(x)$ satisfies (LABEL:eq:equation_for_vpx_neq_vpy_d-2).
 So when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-2})$ is constant for all $b_{d-2}$, denote it by $u_{l}(b_{0},b_{1},\dots,b_{d-3})$.
@@
 Proceeding recursively, we obtain
-$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{z}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},\dots,a_{s},a_{s+1})$ (3.10)
-$+\rho p^{d}e^{u_{l}(b_{0},\dots,b_{s},b_{s+1})}-(p^{d+1+s}+p^{2d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},\dots,b_{s},b_{s+1})=0$
+$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=l\end{subarray}}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},\dots,a_{s},a_{s+1})$
+$+\rho p^{d}e^{u_{l}(b_{0},\dots,b_{s},b_{s+1})}-(p^{d+1+s}+p^{2d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},\dots,b_{s},b_{s+1})=0,$
 $\forall s=0,1,\dots,d-2$.
@@
 If we fix $b_{i}=c_{i}$, $0\leq i\leq s$, $b_{s+1}\neq c_{s+1}$ and suppose $u_{l}(c_{0},\dots,c_{s},b_{s+1},\dots,b_{d-1})$ is independent of $b_{s+2},\dots,b_{d-1}$, denote it by $u_{l}(c_{0},\dots,c_{s},b_{s+1})$.
-Next, we prove that $u_{l}(c_{0},\dots,c_{s},b_{s+1})$ is independent of $b_{s+1}$: From (3.6) and (LABEL:coefficient_of_u(x)), we can get
-$p^{d}Du(x)=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.11)
+Next, we prove that $u_{l}(c_{0},\dots,c_{s},b_{s+1})$ is independent of $b_{s+1}$: From (3.8) and (LABEL:eq:coefficient_of_u(x)), we can get
+$p^{d}Du(x)=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.12)
 $-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
-$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
-$+\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
-we compute
+$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
+$+\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1}).$
+We first compute
 $\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
-$=$ $p^{2(s+1)}(p^{d-s-2}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
-$+(p-1)p^{d-s-3}u_{l}(c_{0},\dots,c_{s+1})+\dots+(p-1)u_{l}(c_{0},\dots,c_{d-2})+u(y))$
+$=$ $p^{2(s+1)}\Biggl(p^{d-s-2}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
+$+(p-1)p^{d-s-3}u_{l}(c_{0},\dots,c_{s+1})+\dots+(p-1)u_{l}(c_{0},\dots,c_{d-2})+u(y)\Biggr)$
 $=$ $p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
@@
 $\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)=$ $\sum_{j=s+2}^{d-1}(p-1)p^{d-1-j}p^{2j}u_{l}(c_{0},\dots,c_{s},b_{s+1})$
-$=$ $(p^{2d-1}-p^{d+s+1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
+$=$ $(p^{2d-1}-p^{d+s+1})u_{l}(c_{0},\dots,c_{s},b_{s+1}).$
 Substituting into (3.12) yields
 $p^{d}Du(x)$ (3.13)
-$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
-$+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)-(p^{d+s+1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
-$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
-$+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
+$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
+$+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
+$-(p^{d+s+1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
+$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
+$+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
+$-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1}).$
+Then equation (LABEL:eq:equation_for_x_neq_y) becomes
+$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
+$+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
+$-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})+\rho p^{d}e^{u_{l}(c_{0},\dots,c_{s},b_{s+1})}=0.$
 So when $p^{2d}\rho\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}$ depend only on $c_{0},c_{1},\dots,c_{s}$, $\forall s=0,1,\dots,d-2$.
 Then we get the result by induction.
-In conclusion, when $v_{p}(x)=l=v_{p}(y)$, $p^{2d}\rho\leq p^{d+1+s}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(c_{0},\dots,c_{s})$ is constant.
+In conclusion, when $v_{p}(x)=l=v_{p}(y)$, $p^{2d}\rho\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(c_{0},\dots,c_{s})$ is constant.
 Then when
-$\rho$ $\leq p^{-2d}\text{min}_{\begin{subarray}{c}-1\leq s\leq d-2\\ 0\leq l\leq m-1\end{subarray}}\{p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}\}$
+$\rho$ $\leq p^{-2d}\min_{\begin{subarray}{c}-1\leq s\leq d-2\\ 0\leq l\leq m-1\end{subarray}}\{p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}\}$
 $=p^{-2d}(p^{d}+p^{d-1}-(1+p^{1-m})p^{d-1})$
 $=p^{-2d}(p^{d}-p^{d-m})$
-$=p^{-d}-p^{-d-m}$
-$u(x)$ is radial symmetric.
-By Lemma 3.4 and Lemma 3.5, we can see that $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, and if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ with $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$, which means that $u(x)$ is radial symmetric.
+$=p^{-d}-p^{-d-m},$
+$u(x)$ is radially symmetric.
+By Lemma 3.4 and Lemma 3.6, we can see that $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, and if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ with $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$.
 ∎
-From our numerical results, we can also see that the radial symmetric solutions have monotonicity. $u(x)$ increases with the distance from the singular point $y$.
+By observing these radially symmetric solutions from numerical experiments, we find that the solution has monotonicity. $u(x)$ increases with the distance from the singular point $y$.
 Theorem 3.8.
-Assume $\rho\leq p^{-d}-p^{-d-m}$ and $v_{p}(x)=l=v_{p}(y)$, then $u_{l}(c_{0},\dots,c_{d-1})<u_{l}(c_{0},\dots,c_{d-2})<\dots<u_{l}(c_{0})<u_{l}$
+Assume $\rho\leq p^{-d}-p^{-d-m}$ and $v_{p}(x)=l=v_{p}(y)$, then the radially symmetric solution satisfies $u_{l}(c_{0},\dots,c_{d-1})<u_{l}(c_{0},\dots,c_{d-2})<\dots<u_{l}(c_{0})<u_{l}$
 Proof.
-Let $v_{p}(z)=k$, by Theorem 3.6, when $\rho\leq p^{-d}-p^{-d-m}$, $u(x)$ is radial symmetric, so the equation variables can be reduced from $m(p-1)p^{d-1}$ to $p+m$.
+From Theorem 3.7, when $\rho\leq p^{-d}-p^{-d-m}$, the solution is radially symmetric, so the equation variables can be reduced from $m(p-1)p^{d-1}$ to $p+m$.
 Then denote these variables by $u_{0},\dots,u_{l-1},u_{l},u_{l}(c_{0}),\dots,u_{l}(c_{0},\dots,c_{d-1}),u_{l+1},\dots,u_{m-1}$.
-When $x\neq y$, equation (3.4) becomes
-$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}+p^{d}\rho e^{u_{l}(x)}=0$ (3.13)
-When $b_{i}=c_{i}$ for $0\leq i\leq s$, then from (LABEL:Du_3), we have
+When $x\neq y$, by letting $v_{p}(z)=k$, equation (3.4) becomes
+$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.14)
+$-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}+p^{d}\rho e^{u_{l}(x)}=0.$
+When $b_{i}=c_{i}$ for $0\leq i\leq s$, then from (LABEL:eq:compute_p^dDu_index_simplified), we have
 $\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s})$ (3.15)
@@
 $=$ $\sum_{j=0}^{s}\sum_{a_{j}\neq c_{j}}p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{d+s}(p-1-\delta_{s,-1})u_{l}(c_{0},\dots,c_{s})$
-$+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
-$=$ $\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
+$+p^{2(s+1)}\left(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
+$-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
+$=$ $\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{2(s+1)}\left(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
 $-(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
@@
 Suppose $u_{l}(c_{0},\dots,c_{s+1})>\dots>u_{l}(c_{0},\dots,c_{d-1})$, now we want to compare $u_{l}(c_{0},\dots,c_{s+1})$ with $u_{l}(c_{0},\dots,c_{s})$.
-First, by (LABEL:eq2_1) and (LABEL:eq2_2), they satisfy the following equations:
-$(p-1)p^{d-1}(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k})+\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.15)
-$+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
+First, by (LABEL:eq:equation_radial_symmetric_x_neq_y) and (LABEL:eq:equation_radial_symmetric_x_neq_y_crutial_part), they satisfy the following equations:
+$(p-1)p^{d-1}\left(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k}\right)+\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.16)
+$+p^{2(s+1)}\left(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
 $-(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s})}=0$
-$(p-1)p^{d-1}(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k})+\sum_{j=0}^{s+1}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.16)
-$+p^{2(s+2)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
-$-(2p^{s+2}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}=0$
-subtracting (LABEL:eq2_4) from (LABEL:eq2_3) gives
+and
+$(p-1)p^{d-1}\left(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k}\right)+\sum_{j=0}^{s+1}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.17)
+$+p^{2(s+2)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
+$-(2p^{s+2}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}=0.$
+Subtracting (LABEL:eq:equation_radial_symmetric_up_to_s+1) from (LABEL:eq:equation_radial_symmetric_up_to_s) gives
 $0=$ $p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}-p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s})}$
@@
 $+(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
-$+(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+(p-1-\delta_{s+1,0})p^{d+s}u_{l}(c_{0},\dots,c_{s})$
+$+(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+(p-1-\delta_{s+1,0})p^{d+s}u_{l}(c_{0},\dots,c_{s}),$
 which is
@@
 $-(2p^{s+2}+p^{s+1}-p^{s}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})$
-$+(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
-$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
+$+(p^{2}-1)p^{2(s+1)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
+$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s}).$
 Then by induction hypothesis and noting that
-$(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}+1)$
-$=$ $(p^{2}-1)p^{2(s+1)}((p-1)p^{d-2}\frac{p^{-s-1}-p^{-d+2}}{p-1}+1)$
-$=$ $p^{d-1}(p^{s+2}-p^{s})$
+$(p^{2}-1)p^{2(s+1)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}+1\right)$
+$=$ $(p^{2}-1)p^{2(s+1)}\left((p-1)p^{d-2}\frac{p^{-s-1}-p^{-d+2}}{p-1}+1\right)$
+$=$ $p^{d-1}(p^{s+2}-p^{s}),$
 we obtain
@@
 $-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})$
-$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
+$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s}),$
 which is
@@
 $-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s+1})$
-$+p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
-since $\rho\leq p^{-d}-p^{-d-m}$, then by the monotonicity of $f_{s}(x)=\rho e^{x}-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})x$, we have $u_{l}(c_{0},\dots,c_{s+1})<u_{l}(c_{0},\dots,c_{s})$.
+$+p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s}).$
+Then by the monotonicity of $f_{s}(x)=\rho e^{x}-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})x$, if $\rho\leq p^{-d}-p^{-d-m}$, we have $u_{l}(c_{0},\dots,c_{s+1})<u_{l}(c_{0},\dots,c_{s})$.
 Remark 3.9.
@@
 $f_{s}(u_{l}(c_{0},\dots,c_{s+1}))-f_{s}(u_{l}(c_{0},\dots,c_{s}))$
-$=$ $(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}((u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{j}))$
-$+(u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{d-1}))$
+$=$ $(p^{2}-1)p^{2(s+1)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}((u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{j})\right)$
+$+(u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{d-1})),$
 so by the monotonicity of the solution and $f_{s}(x)$, the gap between them decreases.
-This theorem provides a sharper estimate of the upper bound for the solution. We now refine Lemma 3.2:
+Remark 3.10.
+There always exists radially symmetric solutions as we can view the $p+d$ points as a finite weighted graph.
+Then from the results in [13] we obtain the existence of the solution, regardless of the value of $\rho$.
+Remark 3.11.
+The upper bound of $\rho$ that ensures $u$ is radially symmetric and monotone is highly dependent on the upper bound of $u$. And Lemma 3.2 yields only a crude estimate for the upper bound of $u$.
+From Theorem 3.8, we can provides a sharper estimate of the upper bound for the solution. We now refine Lemma 3.2:
 Lemma 3.12.
-$u(i)<\begin{cases}\text{ln}\frac{p}{p-2},\ &p>2\\ 4,&p=2\end{cases}$.
+If $\rho\leq p^{-d}-p^{-d-m}$, then the radially symmetric solution $u$ satisfies
+$u(i)<\begin{cases}\ln\left(\frac{p}{p-2}\right),\ &p>2;\\ \ln 4,&p=2.\end{cases}$
 Proof.
@@
 $(p-2)p^{d-1}e^{M}<p^{d}$
-thus $M<\text{ln}\frac{p}{p-2}$.
-When $p=2$, there are at least $2^{d-2}$ such $M$, so $M<4$. ∎
-Using this improved bound, we refine Theorem (3.6) and Theorem (3.7):
+thus $M<\ln\left(\frac{p}{p-2}\right)$.
+When $p=2$, there are at least $2^{d-2}$ such $M$, so $M<\ln 4$. ∎
+Using this improved bound, we can refine Theorem (3.7) and Theorem (3.8) by method of continuity:
 Theorem 3.13.
-When $\rho\leq\begin{cases}\frac{p-2}{p}(1-p^{-m}),\ &p>2\\ \frac{1-2^{-m}}{4},&p=2\end{cases}$, then the solution is radial symmetric and monotone.
+When $\rho\in(0,\rho^{*})$, the radially symmetric solution is monotone with a universal upper bound $M$ independent of $d$ and $\rho$, where $\rho^{*}=\theta_{p,m}e^{-M}$ and $\theta_{p,m}$ is the smallest nonzero eigenvalue of $\boldsymbol{A}$.
+Furthermore, when $\rho\in(0,(1-p^{-m})e^{-M}]$, the solution $u$ must be radially symmetric.
+Proof.
+Denoting
+[…]
+and
+[…]
+First, from Theorem 3.8, we have $(0,p^{-d}-p^{-d-m}]\in S$.
+Next, suppose $\rho_{0}\in S$, and the corresponding radially symmetric solution is $\boldsymbol{u}_{0}$. Then by the same argument of Lemma 3.12, we have
+$u_{0}(i)\leq M<\begin{cases}\ln\left(\frac{p}{p-2}\right),\ &p>2;\\ \ln 4,&p=2,\end{cases}$
+which means any radially symmetric solution have a universal upper bound, so $M$ is well defined. Now considering
+$F:\mathbb{R}\times V_{rad}$ $\to V_{rad}$
+$(\rho,\boldsymbol{u})$ $\mapsto F(\rho,\boldsymbol{u})=\boldsymbol{A}\boldsymbol{u}+\rho e^{\boldsymbol{u}}-\rho p^{d}\boldsymbol{e}_{y},$
+[…]
+$L_{\rho_{0}}:=D_{\boldsymbol{u}}F(\rho_{0},\boldsymbol{u}_{0})=\boldsymbol{A}+\rho_{0}\text{diag}(e^{\boldsymbol{u}_{0}}):V_{rad}\to V_{rad}$
+is nondegenerate since $\rho_{0}e^{u_{0}(i)}<\rho^{*}e^{M}=\theta_{p,m}$, then by Implicit Function Theorem, there exists $\varepsilon>0$ s.t.
+$(\rho_{0}-\varepsilon,\rho_{0}+\varepsilon)$ $\to V_{rad}$
+$\rho$ $\mapsto\boldsymbol{u}_{\rho}$
+satisfies $F(\rho,\boldsymbol{u}_{\rho})=0$ and $\boldsymbol{u}_{\rho_{0}}=\boldsymbol{u}_{0}$.
+Since we can take $\varepsilon$ small enough to ensure $\boldsymbol{u}_{\rho}$ is radially symmetric, so $S$ is open in $(0,\rho^{*})$.
+[…]
+$f_{s}(u_{l}(c_{0},\dots,c_{s+1}))>f_{s}(u_{l}(c_{0},\dots,c_{s})),$
+a contradiction. So $S$ is closed in $(0,\rho^{*})$.
+In conclusion, $S=(0,\rho^{*})$. Then by the same argument of Theorem 3.7, we obtain when $\rho\in(0,(1-p^{-m})e^{-M}]$, the solution $u$ must be radially symmetric. ∎
 3.2 Uniqueness of the solution
 In this part, we discuss the uniqueness of the solution.
-Numerical experiments indicates that for large $\rho$, solutions may not unique, and we can get some solutions that break the properties of radial symmetric.
-Then by Lemma 3.4 and Lemma 3.5, we can construct many non-radial symmetric solutions.
-A natural question is whether uniqueness holds under the conditions of $\rho$ guaranteeing radial symmetric of the solution.
-The answer is affirmative with a slighter stricter condition on $\rho$.
-Uniqueness is straightforward on the finite quotient. If $\boldsymbol{u}$ and $\boldsymbol{v}$ both solve (3.4). Then
-$\boldsymbol{A}(\boldsymbol{u}-\boldsymbol{v})+\rho(e^{\boldsymbol{u}}-e^{\boldsymbol{v}})=\boldsymbol{0}$
-By the mean value theorem applied for each component, there exists $\boldsymbol{\xi}$ such that
-$(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})(\boldsymbol{u}-\boldsymbol{v})=\boldsymbol{0}$
-Lemma 3.9 gives $e^{\xi_{i}}\leq\frac{p}{p-2}$, which is independent of $\rho$.
-Since $\boldsymbol{A}$ is semi-negative definite, if $\rho\frac{p}{p-2}$ does not exceed the first non-zero eigenvalue of $\boldsymbol{A}$, then $(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})$ is invertible, then we get the uniqueness.
-So the next step is to study the eigenvalue of $\boldsymbol{A}$, i.e.
+Numerical experiments indicates that for large $\rho$, solutions may not unique, and we can get some solutions that break the properties of radially symmetric.
+Then by Lemma 3.4 and Lemma 3.6, we can construct many non-radially symmetric solutions.
+Besides, the monotonicity of the solutions may not be ensured for large $\rho$.
+For instance, if we take $p=3$, $m=1$, $d=2$, $\rho=5$ and $y=1$, then there exists non-monotone solution $\boldsymbol{u}_{nmon}$ and non-radially symmetric solution $\boldsymbol{u}_{nrad}$, see in Table 1.
+$\boldsymbol{u}_{nmon}$ $\boldsymbol{u}_{nrad}$
+$u(1)$ -18.2689 -22.2707
+$u(2)$ -1.3831 -0.4886
+$u(4)$ 1.4168 1.9684
+$u(5)$ -1.3831 -0.4886
+$u(7)$ 1.4168 -8.7704
+$u(8)$ -1.3831 -0.4886
+Table 1: Non-monotone solutions and non-radially symmetric solutions
+A natural question is whether uniqueness holds under the conditions of $\rho$ guaranteeing radially symmetric and monotonicity of the solution, the answer is affirmative.
+In order to investigate it, our first step is to study the eigenvalue of $\boldsymbol{A}$, i.e.
 the spectrum of $D$ on the finite quotient $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$.
@@
 Denoting $\boldsymbol{P}\in\mathbb{R}^{m\times m}$ with $(\boldsymbol{P})_{ij}=p^{-|i-j|}$.
-And $\boldsymbol{T}\in\mathbb{R}^{n\times n}$ be the circulant part on diagonal with $(\boldsymbol{T})_{ij}=(\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}-1)(1-\delta_{zx})$ (i.e.
+And $\boldsymbol{T}\in\mathbb{R}^{n\times n}$ be the circulant part on diagonal with $(\boldsymbol{T})_{ij}=\left(\frac{|z|_{p}|x|_{p}}{|z-x|_{p}^{2}}-1\right)(1-\delta_{zx})$ (i.e.
 $\boldsymbol{T}$ is zero on diagonal and $\boldsymbol{T}$ is independent of $v_{p}(z)=v_{p}(x)$).
 Then we can write $\boldsymbol{B}$ as
-$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-\text{diag}((\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn})$
+$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-\text{diag}((\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{1}_{mn}).$
 First, we calculate
@@
 $=$ $\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}\boldsymbol{1}_{n}+\boldsymbol{I}_{m}\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n}$
-$=$ $n\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}+\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n}$
+$=$ $n\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n}+\boldsymbol{1}_{m}\otimes\boldsymbol{T}\boldsymbol{1}_{n},$
 let $t_{0}$ be the row sum of $\boldsymbol{T}$. Then
@@
 $=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}\otimes\boldsymbol{1}_{n})+\text{diag}(\boldsymbol{1}_{m}\otimes t_{0}\boldsymbol{1}_{n})$
-$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn}$
+$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn},$
 hence
-$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn})$
+$\boldsymbol{B}=\boldsymbol{P}\otimes\boldsymbol{1}_{n}\boldsymbol{1}_{n}^{\top}+\boldsymbol{I}_{m}\otimes\boldsymbol{T}-(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}+t_{0}\boldsymbol{I}_{mn}).$
 Let $\boldsymbol{F}$ be the DFT matrix, as any circulant matrix can be diagonalized by a DFT matrix, denoting the eigenvalue of $\boldsymbol{T}$ is $t_{0},t_{1}\dots,t_{n-1}$, which is real and $t_{0}$ is the largest as $\boldsymbol{T}$ is a symmetric circulant matrix and the elements in $\boldsymbol{T}$ are nonnegative.
@@
 $=$ $\boldsymbol{P}\otimes\begin{pmatrix}n&\\ &\boldsymbol{O}_{n-1}\end{pmatrix}$
-$=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &\boldsymbol{O}_{mn-m}\end{pmatrix}$
-$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}\boldsymbol{T}\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda}$
+$=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &\boldsymbol{O}_{mn-m}\end{pmatrix}$,
+$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{I}_{m}\otimes\boldsymbol{T})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}\boldsymbol{T}\boldsymbol{F}=\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda},$
 $\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{F}^{-1}\boldsymbol{I}_{n}\boldsymbol{F}$
-$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}$
-$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(t_{0}\boldsymbol{I}_{mn})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=t_{0}\boldsymbol{I}_{mn}$
+$=$ $n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n},$
+$\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(t_{0}\boldsymbol{I}_{mn})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=t_{0}\boldsymbol{I}_{mn}.$
 Thus
 $\boldsymbol{I}_{m}\otimes\boldsymbol{F}^{-1}(\boldsymbol{B})\boldsymbol{I}_{m}\otimes\boldsymbol{F}=$ $\begin{pmatrix}n\boldsymbol{P}&\\ &0\end{pmatrix}+\boldsymbol{I}_{m}\otimes\boldsymbol{\Lambda}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\otimes\boldsymbol{I}_{n}-t_{0}\boldsymbol{I}_{mn}$
-$=$ $\begin{pmatrix}n\boldsymbol{P}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})&\\ &(t_{i}-t_{0})\boldsymbol{I}_{m}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\end{pmatrix}$
+$=$ $\begin{pmatrix}n\boldsymbol{P}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})&\\ &(t_{i}-t_{0})\boldsymbol{I}_{m}-n\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})\end{pmatrix}$.
 Denoting $s_{j}$ the row sum of $\boldsymbol{P}$, we have
@@
 $=$ $\sum_{i=0}^{j-1}p^{-(j-i)}+\sum_{i=j}^{m-1}p^{-(i-j)}$
-$=$ $\frac{1+p-p^{-j}-p^{j+1-m}}{p-1}$
-So the eigenvalue of $\boldsymbol{B}$ is $n\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))=(p-1)p^{d-1}\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$ and $t_{i}-t_{0}-ns_{j}=t_{i}-t_{0}-p^{d-1}(1+p-p^{-j}-p^{j+1-m})$, where $i=1,\dots,n-1;\ j=0,\dots,m-1$.
-Hence, the eigenvalue of $\boldsymbol{B}$ is $(1-p^{-1})\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$ and $p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}$.
+$=$ $\frac{1+p-p^{-j}-p^{j+1-m}}{p-1}.$
+So the eigenvalue of $\boldsymbol{B}$ is
+$n\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))=(p-1)p^{d-1}\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$
+and
+$t_{i}-t_{0}-ns_{j}=t_{i}-t_{0}-p^{d-1}(1+p-p^{-j}-p^{j+1-m}),$
+where $i=1,\dots,n-1;\ j=0,\dots,m-1$. Hence, the eigenvalue of $\boldsymbol{A}$ is
+$(1-p^{-1})\sigma(\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m}))$
+and
+$p^{-d}(t_{i}-t_{0})-\frac{1+p-p^{-j}-p^{j+1-m}}{p}.$
 Note that $\boldsymbol{P}-\text{diag}(\boldsymbol{P}\boldsymbol{1}_{m})$ is symmetric, then there exists an orthogonal matrix $\boldsymbol{Q}\in\mathbb{R}^{m\times m}$ to diagonalize it.
@@
 Lemma 3.15.
-The smallest non-zero eigenvalue of $\boldsymbol{A}$ is $\begin{cases}(1-p^{-1}),&m=1\\ -(1-p^{-1})\lambda_{1},&m>1\end{cases}$
+The smallest non-zero eigenvalue of $\boldsymbol{A}$ is
+$\theta_{p,m}=\begin{cases}(1-p^{-1}),&m=1;\\ -(1-p^{-1})\lambda_{1},&m>1.\end{cases}$
 Proof.
@@
 As $\mu_{0}=0$ with eigenvector $\boldsymbol{1}_{m}$, so by Courant–Fischer min-max theorem, we have
-$\mu_{1}=\text{min}_{\begin{subarray}{c}\boldsymbol{x}\neq\boldsymbol{0}_{m}\\ \boldsymbol{x}\bot\boldsymbol{1}_{m}\end{subarray}}\frac{\boldsymbol{x}^{\top}\boldsymbol{L}\boldsymbol{x}}{\boldsymbol{x}^{\top}\boldsymbol{x}}$
+$\mu_{1}=\min_{\begin{subarray}{c}\boldsymbol{x}\neq\boldsymbol{0}_{m}\\ \boldsymbol{x}\bot\boldsymbol{1}_{m}\end{subarray}}\frac{\boldsymbol{x}^{\top}\boldsymbol{L}\boldsymbol{x}}{\boldsymbol{x}^{\top}\boldsymbol{x}}.$
 Taking $\boldsymbol{x}=\boldsymbol{e}_{0}-\boldsymbol{e}_{m-1}$, we have
@@
 $\leq$ $1$
-with equality if and only if $p=2$.
+with equality if and only if $p=2$, $\forall m$ and $m=1$, $\forall p$.
 Therefore, the smallest nonzero eigenvalue of $\boldsymbol{A}$ is
-$\text{min}_{\begin{subarray}{c}1\leq k\leq m-1\\ 1\leq i\leq d-1\\ 0\leq j\leq m-1\end{subarray}}\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\}$
-$=$ $\text{min}\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
-$=$ $\begin{cases}(1-p^{-1}),&m=1\\ -(1-p^{-1})\lambda_{1},&m>1\end{cases}$
+$\theta_{p,m}=$ $\min_{\begin{subarray}{c}1\leq k\leq m-1\\ 1\leq i\leq d-1\\ 0\leq j\leq m-1\end{subarray}}\left\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\right\}$
+$=$ $\min\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
+$=$ $\begin{cases}(1-p^{-1}),&m=1;\\ -(1-p^{-1})\lambda_{1},&m>1.\end{cases}$
 Theorem 3.16.
-When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then the solution of (3.2) is unique.
+When $\rho\in(0,\rho^{*})$, then the solution of (3.2) is unique.
 Proof.
-By Lemma 3.9, when $p>2$, we have
-$\rho\frac{p}{p-2}\leq(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$
-thus $\rho\leq(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$.
-When $p=2$,we have
-$4\rho\leq\frac{1}{2}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$
-thus $\rho\leq\frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1})$. ∎
+As $\theta_{p,m}\leq(1-p^{-m})$ by Lemma 3.15, then the solution must be radially symmetric. If $\boldsymbol{u}$ and $\boldsymbol{v}$ both solve (3.4). Then
+$\boldsymbol{A}(\boldsymbol{u}-\boldsymbol{v})+\rho(e^{\boldsymbol{u}}-e^{\boldsymbol{v}})=\boldsymbol{0}.$
+By the mean value theorem applied for each component, there exists $\boldsymbol{\xi}$ such that
+$(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})(\boldsymbol{u}-\boldsymbol{v})=\boldsymbol{0}.$
+Theorem 3.13 gives $e^{\xi_{i}}\leq e^{M}$, which is independent of $\rho$ and $d$. Besides, as $\boldsymbol{A}$ is semi-negative definite and
+$\rho e^{\xi_{i}}<\rho^{*}e^{M}=\theta_{p,m},$
+then $\rho e^{\xi_{i}}$ does not exceed the first non-zero eigenvalue of $\boldsymbol{A}$. Then $(\boldsymbol{A}+\rho\text{diag}e^{\boldsymbol{\xi}})$ is invertible, then we get the uniqueness. ∎
+Remark 3.17.
+The order of $\rho^{*}$ is optimal. Since $\sum_{i}e^{u_{i}}=p^{d}$, we have $m(p-1)p^{d-1}e^{M}\geq p^{d}$, therefore $M>\ln\left(\frac{p}{m(p-1)}\right)$, which means
+$\rho^{*}<\frac{m(p-1)}{p}(1-p^{-m}).$
+On the other hand, since
+$M<\begin{cases}\ln\left(\frac{p}{p-2}\right),\ &p>2;\\ \ln 4,&p=2,\end{cases}$
+which means
+$\rho^{*}>\begin{cases}\frac{1}{4}\theta_{2,m},&p=2;\\ (1-2p^{-1})\theta_{p,m},\ &p>2.\end{cases}=\begin{cases}\frac{1}{8},&p=2,\ m=1;\\ -\frac{1}{8}\lambda_{1},&p=2,\ m>1;\\ (1-2p^{-1})(1-p^{-1}),\ &p>2,\ m=1;\\ -(1-2p^{-1})(1-p^{-1})\lambda_{1},&p>2,\ m>1.\end{cases}$
 4 Mean field equation on the Tate curve
-Having studied the solution on finite quotient, we now consider the solution on the Tate curve.
 Since the finite quotient converges to the Tate curve when $d\rightarrow\infty$, we investigate whether the solution converge accordingly.
-The answer is affirmative.
+From the numerical experiments, we observe that the solution on the finite quotient is convergent.
 Moreover, the limiting solution retains properties like uniqueness and monotonicity.
 4.1 Convergence to the Tate curve
-In this part, we will show that the solution on the finite quotient will converge to the solution on the Tate curve, thus we can get the existence of the solution.
-From the numerical experiments, we can indeed observe that the solution on the finite quotient is convergent.
-Next, we will prove it rigorously.
 Theorem 4.1.
-When $\rho\leq\begin{cases}\frac{p-2}{p}(1-p^{-m}),\ &p>2\\ \frac{1-2^{-m}}{4},&p=2\end{cases}$, then there exists a solution $u$ satisfying (3.1).
+When $\rho\in(0,\rho^{*})$, then there exists a solution $u$ satisfying (3.1).
 Proof.
-For convenience, we denote $X_{m}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with the Haar measure $d\mu=\frac{dx}{|x|_{p}}$ and $X_{m,d}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ with the Haar measure $d\mu_{d}=p^{-d}d\delta$.
-Clearly, $\mu_{d}\overset{*}{\rightharpoonup}\mu$ as $d\rightarrow\infty$.
+For convenience, we denote $X_{m}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$ with the Haar measure $d\mu=\frac{dx}{|x|_{p}}$ and $X_{m,d}=\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}/(1+p^{d}\mathbb{Z}_{p})$ with the Haar measure $d\mu_{d}=p^{-d}d\delta$, which can be extended to the measure $\widetilde{\mu_{d}}$ on $X_{m}$.
+Clearly, $\widetilde{\mu_{d}}\overset{*}{\rightharpoonup}\mu$ as $d\rightarrow\infty$.
 Let $D_{d}$ denote the operator on the finite quotient.
@@
 with
-$\int_{X_{m,d}}e^{u_{d}}d\mu_{d}=1$ (4.2)
+$\int_{X_{m,d}}e^{u_{d}}d\mu_{d}=1.$ (4.2)
 Since $D_{d}u_{d}=\rho p^{d}e_{y}-\rho e^{u_{d}}$, we have $|D_{d}u_{d}|\leq\rho p^{d}e_{y}+\rho e^{u_{d}}$, then
-$\int_{X_{m,d}}|D_{d}u_{d}|d\mu_{d}\leq 2\rho$ (4.3)
+$\int_{X_{m,d}}|D_{d}u_{d}|d\mu_{d}\leq 2\rho.$ (4.3)
 Define measures $\nu_{d}=e^{u_{d}}\mu_{d}$ and $\omega_{d}=D_{d}u_{d}\mu_{d}$ on $X_{m,d}$.
@@
 $\widetilde{\nu_{d}}|_{X_{m,d}}=\nu_{d}$ and $\widetilde{\omega_{d}}|_{X_{m,d}}=\omega_{d}$.
-Since $X_{m}$ is compact, we have $C(X_{m})^{*}=BV(X_{m})$, so by the Banach-Alaoglu theorem, $BV(X_{m})$ is weak-* compact.
-Then from (4.2) and (4.3), we have $\widetilde{\nu_{d}}||_{BV}=1$ and $\widetilde{\omega_{d}}||_{BV}\leq 2\rho$, so there exists $\nu$, $\omega\in BV(X_{m})$ and a subsequence $d_{k}$ such that
+On the other hand, the function on $X_{m,d}$ can extend to a locally constant function on $X_{m}$.
+Since $X_{m}$ is compact, we have $C(X_{m})^{*}=\mathscr{M}(X_{m})$, so by the Banach-Alaoglu theorem, $\mathscr{M}(X_{m})$ is weak-* compact.
+Then from (4.2) and (4.3), we have $\widetilde{\nu_{d}}||_{\mathscr{M}}=1$ and $\widetilde{\omega_{d}}||_{\mathscr{M}}\leq 2\rho$, so there exists $\nu$, $\omega\in\mathscr{M}(X_{m})$ and a subsequence $d_{k}$ such that
 $\widetilde{\nu_{d_{k}}}\overset{*}{\rightharpoonup}\nu,\ \widetilde{\omega_{d_{k}}}\overset{*}{\rightharpoonup}\omega$
-as $\nu_{d}<<\mu_{d}$ and $\frac{d\nu_{d}}{d\mu_{d}}=e^{u_{d}}\leq C_{p}$ uniformly in $d$ by Lemma 3.9, we have $\nu<<\mu$, denoting $\nu=e^{\widetilde{u}}\mu$ for some function $\widetilde{u}$.
-Take any locally constant function $\phi$ on $X_{m}$, so when $d$ is large enough, $\phi$ is well-defined on $X_{m,d}$ as $X_{m}$ is compact.
+as $\nu_{d}<<\mu_{d}$ and $\frac{d\nu_{d}}{d\mu_{d}}=e^{u_{d}}\leq e^{M}$ uniformly in $d$ by Theorem 3.13, we have $\nu<<\mu$, denoting $\nu=e^{\bar{u}}\mu$ for some function $\bar{u}$.
+Taking any locally constant function $\phi$ on $X_{m}$, when $d$ is large enough, $\phi$ is well-defined on $X_{m,d}$ as $X_{m}$ is compact.
 Multiply (4.1) by $\phi$ and integrate over $X_{m,d}$, we have
-$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}+\rho\int_{X_{m,d}}e^{u_{d}}\phi d\mu_{d}=\rho\phi(y)$
+$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}+\rho\int_{X_{m,d}}e^{u_{d}}\phi d\mu_{d}=\rho\phi(y),$
+and it can be rewritten as
+$\int_{X_{m}}\phi d\widetilde{\omega_{d}}+\rho\int_{X_{m}}\phi d\widetilde{\mu_{d}}=\rho\phi(y).$
 By taking a limit of the subsequence $d_{k}$, we get
-$\int_{X_{m}}\phi d\omega+\rho\int_{X_{m}}\phi d\nu=\rho\phi(y)$ (4.4)
+$\int_{X_{m}}\phi d\omega+\rho\int_{X_{m}}\phi d\nu=\rho\phi(y).$ (4.4)
 Decompose $u_{d}=v_{d}+\rho G_{d}$, where $G_{d}$ is the Green’s function on finite quotient, then $v_{d}$ satisfies
-$D_{d}v_{d}=-\rho e^{u_{d}}+\frac{\rho}{V}=-\rho e^{\rho G_{d}}e^{v_{d}}+\frac{\rho}{V}$
+$D_{d}v_{d}=-\rho e^{u_{d}}+\frac{\rho}{V}=-\rho e^{\rho G_{d}}e^{v_{d}}+\frac{\rho}{V},$
 so
-$v_{d}(x)=\int_{X_{m,d}}G_{d}(x,y)(-\rho e^{u_{d}(y)}+\frac{\rho}{V})d\mu_{d}(y)$
-Since $G(x,y)\sim\frac{p(p-1)}{p+1}\text{log}_{p}d(x,y)$ as $d(x,y)\rightarrow 0$ [14], and $X_{m}$ is compact, then
+$v_{d}(x)=\int_{X_{m,d}}G_{d}(x,y)\left(-\rho e^{u_{d}(y)}+\frac{\rho}{V}\right)d\mu_{d}(y).$
 […]
-and we have $\int_{B_{\epsilon}(x)}|G(x,y)|d\mu(y)\leq C\epsilon|\text{ln}\epsilon|$, so $G\in L^{1}(X_{m})$. and we notice that $\rho e^{u_{d}(y)}+\frac{\rho}{V}$ is bounded, we have
 […]
-so $v_{d}\in L^{\infty}(X_{m,d})$.
+And we have
+$\int_{B_{\epsilon}(x)}\text{log}_{p}d(x,y)d\mu(y)=$ $\int_{x+p^{v_{p}(x)+N}\mathbb{Z}_{p}}\text{log}_{p}d(x,y)d\mu(y)$
+$=$ $\int_{\mathbb{Z}_{p}}\text{log}_{p}(p^{-N}|t|_{p})p^{-N}dt$
+$=$ $-\left(N+\frac{1}{p-1}\right)p^{-N}$
+and
+$\int_{B_{\epsilon}(x)}d(x,y)^{n}d\mu(y)=$ $\int_{x+p^{v_{p}(x)+N}\mathbb{Z}_{p}}d(x,y)^{n}d\mu(y)$
+$=$ $\int_{\mathbb{Z}_{p}}(p^{-N}|t|_{p})^{n}p^{-N}dt$
+$=$ $\frac{p^{n}(p-1)}{p^{n+1}-1}p^{-N(n+1)}.$
+So we have
+$\int_{B_{\epsilon}(x)}|G(x,y)|d\mu(y)\leq CNp^{-N}\leq C^{\prime}\epsilon|\ln\epsilon|,$
+thus $G\in L^{1}(X_{m})$. Additionally, as $\rho e^{u_{d}(y)}+\frac{\rho}{V}$ is bounded and $|\widetilde{G_{d}}(\cdot,y)|\leq|G(\cdot,y)|$ [14], we have
+[…]
+so $v_{d}\in L^{\infty}(X_{m,d})$, thus $\widetilde{v_{d}}\in L^{\infty}(X_{m})$.
 Using the self-adjointness of $D$, we have
-$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}=$ $\int_{X_{m,d}}(D_{d}\phi)u_{d}d\mu_{d}$
+$\int_{X_{m,d}}(D_{d}u_{d})\phi d\mu_{d}=$ $\int_{X_{m}}\widetilde{D_{d}u_{d}}\phi d\widetilde{\mu_{d}}$
+$=$ $\int_{X_{m}}(D_{d}\widetilde{u_{d}})\phi d\widetilde{\mu_{d}}$
+$=$ $\int_{X_{m}}(D_{d}\phi)\widetilde{u_{d}}d\widetilde{\mu_{d}}$
 […]
-$=$ $\mathrm{I}+\mathrm{II}$
+$=$ $\mathrm{I}+\mathrm{II},$
 where
-$|\mathrm{I}|=$ $\left|\int_{B_{\epsilon}(y)}(D_{d}\phi)u_{d}d\mu_{d}\right|$
-$\leq$ $\rho\int_{B_{\epsilon}(y)}|D_{d}\phi(x)||G_{d}(x,y)|d\mu_{d}(x)+\int_{B_{\epsilon}(y)}|D_{d}\phi||v_{d}|d\mu_{d}$
-$\leq$ $C_{1}\epsilon|\text{ln}\epsilon|+C_{2}\epsilon$
+[…]
+$\leq$ $\rho\int_{B_{\epsilon}(y)}|D_{d}\phi(x)||\widetilde{G_{d}}(x,y)|d\widetilde{\mu_{d}}(x)+\int_{B_{\epsilon}(y)}|D_{d}\phi||\widetilde{v_{d}}|d\widetilde{\mu_{d}}$
+$\leq$ $C_{1}\epsilon|\ln\epsilon|+C_{2}\epsilon$
 and we have
-$\mathrm{II}=\int_{X_{m,d}\backslash B_{\epsilon}(y)}(D_{d}\phi)u_{d}d\mu_{d}=\int_{X_{m,d}\backslash B_{\epsilon}(y)}(D\phi)u_{d}d\mu_{d}$
-for sufficiently large $d$. And in $X_{m,d}\backslash B_{\epsilon}(y)$, $|u_{d}|\leq C_{\epsilon}$, so there exists a subsequence $u_{d_{k}}\rightarrow u$ such that
-$\mathrm{II}\rightarrow\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\mu$
-letting $\epsilon\rightarrow 0$ we obtain
-$\int_{X_{m,d_{k}}}(D_{d_{k}}u_{d_{k}})\phi d\mu_{d_{k}}\rightarrow\int_{X_{m}}(D\phi)ud\mu=\int_{X_{m}}(Du)\phi d\mu$
+$\mathrm{II}=\int_{X_{m}\backslash B_{\epsilon}(y)}(D_{d}\phi)\widetilde{u_{d}}d\widetilde{\mu_{d}}=\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)\widetilde{u_{d}}d\widetilde{\mu_{d}}$
+for sufficiently large $d$.
+And in $X_{m}\backslash B_{\epsilon}(y)$, since $\widetilde{u_{d}}(x)$ depend only on $|x|_{p}$, $|y|_{p}$ and $|x-y|_{p}$, so $\forall\ x_{0}\in X_{m}\backslash B_{\epsilon}(y)$, $\exists\ r_{0}(x_{0})<\min\{|x_{0}|_{p},|x_{0}-y|_{p}\}$, s.t.
+$\forall\ x_{0}\in B(x_{0},r_{0}(x_{0}))$, we have $|x|_{p}=|x_{0}|_{p}$, $|x-y|_{p}=|x_{0}-y|_{p}$.
+So $\widetilde{u_{d}}(x)$ is locally constant on $B(x_{0},r_{0}(x_{0}))$.
+Since $X_{m}\backslash B_{\epsilon}(y)$ is compact, so $X_{m}\backslash B_{\epsilon}(y)=\bigcup_{i=1}^{K}B(x_{i},r_{i}(x_{i}))$.
+Taking $\delta=\min\{r_{1}(x_{1}),\dots,r_{K}(x_{K})\}$, then $\forall\ x,z\in X_{m}\backslash B_{\epsilon}(y)$ with $|x-z|_{p}<\delta$, we have $\widetilde{u_{d}}(x)=\widetilde{u_{d}}(z)$ as balls on $p$-adic field are either disjoint or concentric.
+And we notice that $|\widetilde{u_{d}}|\leq C_{\epsilon}$, so there exists a subsequence $\widetilde{u_{d_{k}}}$ converges to a locally constant function $u$ on $X_{m}\backslash B_{\epsilon}(y)$ such that
+$\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)\widetilde{u_{d_{k}}}d\widetilde{\mu_{d_{k}}}=\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)(\widetilde{u_{d_{k}}}-u)d\widetilde{\mu_{d_{k}}}+\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\widetilde{\mu_{d_{k}}}$.
+The first term can be controlled by $\widetilde{u_{d_{k}}}-u||_{L^{\infty}(X_{m}\backslash B_{\epsilon}(y))}$, so
+$\mathrm{II}\rightarrow\int_{X_{m}\backslash B_{\epsilon}(y)}(D\phi)ud\mu.$
+Letting $\epsilon\rightarrow 0$ and by the standard diagonal argument, we obtain a convergent function $u$, so
+$\int_{X_{m}}\widetilde{D_{d_{k}}u_{d_{k}}}\phi d\widetilde{\mu_{d_{k}}}\rightarrow\int_{X_{m}}(D\phi)ud\mu=\int_{X_{m}}(Du)\phi d\mu$
 as $k\rightarrow\infty$. Hence $\omega=Du\mu$ in the sense of distributions.
-What’s more,
-$\int_{X_{m}}\phi d\nu=\lim_{k\to\infty}\int_{X_{m,d_{k}}}\phi e^{u_{d_{k}}}d\mu_{d_{k}}=\int_{X_{m}}\phi e^{u}d\mu$
-and by definition,
-$\int_{X_{m}}\phi d\nu=\int_{X_{m}}\phi e^{\widetilde{u}}d\mu$
-so we have $\widetilde{u}=u$ for a.e. $x$. Then by (4.4) we get
-$\int_{X_{m}}(Du)\phi d\mu+\rho\int_{X_{m}}e^{u}\phi d\mu=\rho\phi(y)$
+What’s more, since
+$\int_{X_{m}}\phi d\nu=\lim_{k\to\infty}\int_{X_{m}}\phi e^{\widetilde{u_{d_{k}}}}d\widetilde{\mu_{d_{k}}}=\int_{X_{m}}\phi e^{u}d\mu,$
+and by definition, we have
+$\int_{X_{m}}\phi d\nu=\int_{X_{m}}\phi e^{\bar{u}}d\mu,$
+so $\bar{u}=u$ for a.e. $x$. Then by (4.4) we get
+$\int_{X_{m}}(Du)\phi d\mu+\rho\int_{X_{m}}e^{u}\phi d\mu=\rho\phi(y).$
 […]
 Since the solution we found converges from the solution on finite quotient, it inherits the properties of the solution on finite quotient.
-So the solution $u$ has those radial symmetric properties, what’s more, $u$ is monotone with respect to $d(x,y)$ at the same connected component and $u\leq C_{p}$.
-When $\rho$ breaks the bound of Theorem 4.1, then there may be non-radial symmetric solutions, so that $u$ will not have the universal upper bound, this may prevent the convergence of solutions on the finite quotient, and hence the existence cannot be guaranteed.
+So the solution $u$ has those radially symmetric and monotone properties with a universal upper bound $M$.
+When $\rho\geq\rho^{*}$, then there may exist non-monotone radially symmetric solutions, let alone a non-radially symmetric one, so that $u$ will not have the universal upper bound, this may prevent the convergence of solutions on the finite quotient, and hence the existence cannot be guaranteed.
 4.2 Uniqueness of the solution
-Similarly from the argument of the solution in finite quotient, we now address uniqueness on the Tate curve under the same constraints of $\rho$.
-Although the solution on finite quotient is unique when $\rho$ is not too large, we can’t ensure the solution on the Tate curve is unique as there may has another solution on the Tate curve but vanishes on finite quotient.
-Actually, the solution is indeed unique under the same constraints of $\rho$.
-We will prove it by a standard method [18].
-First, we show that the solution is unique when $\rho$ is very small, next we prove the linearized solution is non-degenerate under the same constraints of $\rho$, then we can get the uniqueness.
+Although the solution on finite quotient is unique when $\rho\in(0,\rho^{*})$, we can’t ensure uniqueness on the Tate curve as there may has another solution on the Tate curve but vanishes on finite quotient.
+We now address uniqueness on the Tate curve by a standard method [20].
+First, we show that the solution is unique when $\rho\leq\varepsilon$.
+Next, we prove the linearized solution is non-degenerate under the same constraints of $\rho$, then we can get the uniqueness.
 Lemma 4.2.
@@
 Define
-$\phi_{n}=\frac{u_{n}-v_{n}}{\|u_{n}-v_{n}\|_{L^{\infty}}}$
+$\phi_{n}=\frac{u_{n}-v_{n}}{\|u_{n}-v_{n}\|_{L^{\infty}}}.$
 Then $\phi_{n}$ satisfies
-$D\phi_{n}+V_{n}(x)\phi_{n}=0$
+$D\phi_{n}+V_{n}(x)\phi_{n}=0,$
 where $V_{n}(x)=\rho_{n}\frac{e^{u_{n}}-e^{v_{n}}}{u_{n}-v_{n}}$.
-Since $u(x)$ has a universal upper bound by Lemma 3.9, $V_{n}(x)$ converges to 0.
+Since $u(x)\leq M$ by Theorem 3.13, $V_{n}(x)$ converges to 0.
 And it is easy to see that $\phi_{n}$ uniformly converges to the eigenfunction of $D$ with the eigenvalue 0, which is the harmonic function on $\mathbb{Q}_{p}^{\times}/p^{m\mathbb{Z}}$.
@@
 Next, we consider the linearized equation of (3.1) as
-$Df+\rho e^{u}f=0$ (4.5)
-If (4.5) is nondegenerate for $\rho\leq C$, then Lemma 4.2 implies the uniqueness of the solution for such $\rho$. To address this question, we give the following Lemma.
+$Df+\rho e^{u}f=0.$ (4.5)
+If (4.5) is nondegenerate for $\rho\leq\rho^{*}$, then Lemma 4.2 implies the uniqueness of the solution for such $\rho$. To address it, we give the following Lemma:
 Lemma 4.3.
@@
 Suppose that the eigenvalue of LV(x)L-V(x) is $\mu_{1}\leq\mu_{2}\leq\cdots$. Then we have operator inequalities
-$L-\lambda_{k+1}<L-V(x)<L-\lambda_{k}$
+$L-\lambda_{k+1}<L-V(x)<L-\lambda_{k}.$
 Then we have
@@
 for all $i$. Taking $i=k$ and $i=k+1$, we get
-$\mu_{k}<0<\mu_{k+1}$
-which means 0 is not the eigenvalue of LV(x)L-V(x). ∎
-By Lemma 3.9, we can see that $0<\rho e^{u}<\rho\frac{p}{p-2}$.
-And from Lemma 4.3, we can see that if $\rho\frac{p}{p-2}$ is smaller than the first nonzero eigenvalue of D-D, then we can get the uniqueness, similarly, the smallest nonzero eigenvalue of D-D is
-$\text{min}_{\begin{subarray}{c}1\leq k\leq m-1\\ i\geq 1\\ 0\leq j\leq m-1\end{subarray}}\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\}$ (4.6)
-$=$ $\text{min}\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
+$\mu_{k}<0<\mu_{k+1},$
+which means 0 is not the eigenvalue of LV(x)L-V(x). Thus the operator LV(x)L-V(x) is nondegenerate. ∎
+Theorem 4.4.
+When $\rho\in(0,\rho^{*})$, then the solution of (3.1) is unique.
+Proof.
+By Theorem 3.13, we can see that $0<\rho e^{u}<\rho^{*}e^{M}=\theta_{p,m}$. And from Theorem 2.4 and Lemma 3.15, the smallest nonzero eigenvalue of D-D is
+$\min_{\begin{subarray}{c}1\leq k\leq m-1\\ i\geq 1\\ 0\leq j\leq m-1\end{subarray}}\left\{-(1-p^{-1})\lambda_{k},\ p^{i-1}+p^{i-2}-\frac{p^{-j}+p^{j+1-m}}{p}\right\}$ (4.6)
+$=$ $\min\{-(1-p^{-1})\lambda_{1},1-p^{-m}\}$
 […]
-then we have
-Theorem 4.4.
-When $\rho\leq\begin{cases}(1-2p^{-1})(1-p^{-1})(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),\ &p>2\\ \frac{1}{8}(\delta_{m,1}(1+\lambda_{1})-\lambda_{1}),&p=2\end{cases}$, then the solution of (3.1) is unique.
+$=$ $\theta_{p,m}.$
+Then from Lemma 4.3, the linearized operator is nondegenerate, then we obtain the uniqueness. ∎
 Acknowledgement